### v1

# A Neurosymbolic Framework for Interpretable Cognitive Attack Detection in Augmented Reality

Abstract
Augmented Reality (AR) enriches perception by overlaying virtual elements on the physical world. Due to its growing popularity, cognitive attacks that alter AR content to manipulate users’ semantic perception have received increasing attention. Existing detection methods often focus on visual changes, which are restricted to pixel- or image-level processing and lack semantic reasoning capabilities, or they rely on pre-trained vision–language models (VLMs), which function as black-box approaches with limited interpretability. In this paper, we present CADAR, a novel neurosymbolic approach for cognitive attack detection in AR. It fuses multimodal vision–language inputs using neural VLMs to obtain a symbolic perception-graph representation, incorporating prior knowledge, salience weighting, and temporal correlations. The model then enables particle-filter based statistical reasoning–a sequential Monte Carlo method–to detect cognitive attacks. Thus, CADAR inherits the adaptability of pre-trained VLM and the interpretability and reasoning rigor of particle filtering. Experiments on an extended AR cognitive attack dataset show accuracy improvements of up to $10.7\%$ over strong baselines on challenging AR attack scenarios, underscoring the promise of neurosymbolic methods for effective and interpretable cognitive attack detection.
1 Introduction
Augmented Reality (AR) and Mixed Reality (MR) technologies have rapidly advanced in recent years, offering users an immersive perceptual experience by overlaying contextually relevant information and virtual objects onto their physical environment (Zheng and Yuan 2023). As AR/MR are rapidly changing how users perceive, interpret, and interact with the world around them, new vulnerabilities have emerged and attracted increasing research attention. In particular, the reliance on visual and spatial cues in AR/MR creates an implicit trust in the accuracy and authenticity of overlaid content. This trust can be exploited by adversaries to launch cognitive attacks, which use malicious manipulations of virtual content to mislead, distract, or coerce users without their awareness (Xiu, Scargill, and Gorlatova 2025; Cheng et al. 2023; Teymourian et al. 2025). Unlike traditional cybersecurity threats that target data or system integrity, cognitive attacks in AR/MR (such as the examples shown in Fig. 1) operate at the perceptual level, manipulating users’ semantic interpretation through subtle or deceptive visual alterations (Xiu and Gorlatova 2025).
Figure 1: Illustrative AR/MR cognitive attacks. Top row: original scenes; bottom row: manipulated scenes. (a) Text Modification Attack. (b) Visual Modification attack. (c) Removal Attack. (d) Addition Attack.
Prior work on detecting such cognitive attacks often employs computer vision techniques to identify visual changes or adversarial distortions (Qiu et al. 2019; Duan et al. 2022a, b; Li et al. 2024; Acharya et al. 2022). However, these methods rely on pixel/image processing, lacking the ability to model the human perception process that not only involves prior knowledge and contextual information, but also selectively focuses on aspects of the visual field for semantical interpretation. As a result, these methods may struggle to capture smart cognitive attacks manipulating only a small number of pixels and also could suffer from false positives due to benign visual perturbations that do not affect users’ semantic interpretation (Li et al. 2024). More recent studies propose pre-trained vision–language models (VLMs) (Xiu, Scargill, and Gorlatova 2025; Roy et al. 2025) to achieve semantic-level detection, by processing visual features and incorporating human-like knowledge. But the reasoning of pre-trained VLMs is a black-box approach offering limited interpretability. It is hard to leverage domain-specific knowledge or model scene/context evolutions or temporal correlations in cognitive attack detection.
In this paper, we consider cognitive attacks that aim to manipulate users’ semantic perception in AR/MR. We propose CADAR (Cognitive Attack Detection in Augmented Reality), a neuro-symbolic approach encompassing two key components working in concert: (i) a symbolic perception-graph model, which is extracted using pre-trained neural VLM models to represent scene/context evolution at semantic-level, ; and (ii) a particle-filter model, which is commonly used in non-linear and stochastic systems. It applies a sequential Monte Carlo method to the perception graph for reasoning and detecting cognitive attacks, stabilizes inherently noisy or inconsistent VLM outputs, and mitigates errors or missed detections through probabilistic fusion. It also offers granular, flexible temporal detection: by using reference sequences, frame-wise inference localizes attacks per frame, to produce a variable number of labels when needed. Thus, the CADAR approach inherits the adaptability of pre-trained VLM (neural part), as well as the interpretability and reasoning rigor of particle filtering (symbolic part). It semantically interprets and models the sequential scene changes for robust cognitive attacks detection in AR/MR, overcoming limitations in existing work.
Our key contributions are as follows:
  • Neuro-symbolic modeling and reasoning. We present a neuro-symbolic framework tailored to AR/MR cognitive attack detection. A pre-trained neural VLM generates symbolic perception graphs that stores objects, attributes, and relations for explicit, interpretable reasoning.
  • Perception graph representation. We propose a perception graph schema that formally encodes multi-modal scene/context evolution at the perceptual level, integrating outputs from complementary modules (image and text encoders, object detection, and segmentation) for cognitive attack inference.
  • Particle-filter-based statistical reasoning. We are the first to apply particle filtering to symbolic perception graphs. Not only does our method improve the robustness of stochastic VLM outputs, mitigating noise and hallucination, but it also enables interpretable statistical procedures to reason about cognitive attacks.
  • We create the Extended AR-VIM dataset, which is the first publicly available AR cognitive attack dataset, by expanding the existing AR-VIM dataset (Xiu and Gorlatova 2025). CADAR evaluation shows it significantly outperforms several baselines, achieving an overall accuracy of 80.1% compared to 69.4% for the next-best model.
The rest of the paper is organized as follows: Section 2 reviews related work; Section 3 presents our method; Section 4 reports experimental results and Section 5 concludes the paper.
2 Related Work
Cognitive Attacks in AR/MR.
One well-studied category of visual attacks is adversarial attacks, which introduce imperceptible yet malicious perturbations to input data to mislead machine learning models (Zhang et al. 2024a; Mumcu and Yilmaz 2024). In contrast, cognitive attacks target perception-level semantics and aim to alter a human user’s understanding of a scene (Cheng et al. 2023; Xiu and Gorlatova 2025). Augmented Reality (AR) systems, which overlay virtual content onto the physical world to enhance perception and decision-making, are particularly vulnerable to such attacks. However, research on cognitive attacks in AR remains limited. Existing methods often rely on access to the original, unaltered scene for comparison or employ pre-trained VLMs with limited interpretability.
Knowledge Graphs.
Graph-based representations offer a compact and expressive way to encode entities and the possibly complex relations among them, making graphs well-suited for downstream prediction and retrieval tasks. A knowledge graph represents each entity as a node and each semantic relation as an edge, thereby capturing structured world knowledge at scale (Hogan et al. 2021). A scene graph extends this idea to visual understanding: it organizes objects, their attributes, and inter-object relationships into a graph that supports higher-level reasoning over images and videos (Chang et al. 2021). Numerous task-specific graph formulations have also emerged, for example, mission-specific knowledge graphs for situational awareness (Yun et al. 2025), topological graphs for robot navigation (Shah et al. 2022), and interaction networks that model physical relations among objects (Battaglia et al. 2016).
Neuro-symbolic Methods.
Neuro-symbolic approaches couple deep neural networks with symbolic reasoning–i.e., reasoning over explicitly represented knowledge in formal languages, to capitalize on the complementary strengths of subsymbolic pattern recognition and symbolic compositionality (Bhuyan et al. 2024). Within knowledge-graph research, most neuro-symbolic work addresses link prediction or graph completion, where the hybrid models offer interpretable rules, guided learning, and more transparent decision pathways (DeLong, Mir, and Fleuriot 2024). Yet, like standard supervised techniques, many of these solutions still demand sizable labeled datasets. In contrast, our framework introduces a neuro-symbolic reasoning module that combines explicitly encoded rules with statistics, achieving high interpretability while remaining data-efficient.
Particle Filter and Statistical Reasoning.
Particle filters–sequential Monte Carlo algorithms for recursive Bayesian estimation–propagate a set of weighted samples to approximate the posterior state distribution, outperforming classical Kalman filters in strongly non-linear or non-Gaussian settings (Künsch 2013). They are widely used in surveillance, robotics, and navigation, leveraging visual, geometric, and motion cues such as color, texture, and shape (Awal et al. 2023). In CADAR, we model graph nodes and edges as stochastic particles, enabling robust reasoning under noisy or inconsistent VLM outputs.
3 Methodology
In this section, we define the four types of cognitive attacks in AR scenes and introduce our neuro-symbolic framework for detecting such attacks. As illustrated in Fig. 2, the framework consists of two main components: the perception graph model and the particle filter-based detection module. The perception graph encodes each frame as a 3D spatiotemporal structure, where nodes represent objects and edges capture their relationships across time (Fig. 3). These structured graphs are then passed into the attack detection algorithm (Fig. 4), which reasons over the particle filter to identify cognitive attacks.
Figure 2: CADAR system overview. Sequential video frames are converted into temporal perception graphs, which the attack-detection module then analyzes to identify, classify, and localize adversarial attacks.
3.1 Cognitive Attacks
We study four representative adversarial attack types that compromise an AR scene’s semantics (Fig. 1).
Text-modification Attack.
Altering on-scene text, e.g., overlaying “Free Parking” atop a “No Parking” sign, reverses instructions or injects false information, misleading both users and perception systems.
Visual-modification Attack.
Altering an object’s appearance or position (green light to red, stop sign moved to mid-junction) distorts recognition and user judgment.
Removal Attack.
Occluding or deleting safety-critical cues (e.g., an exit sign) erases vital information and breaks expected graph relations.
Addition Attack.
Inserting fictitious objects or labels (e.g., a fake hazard symbol) plants misleading cues that divert attention and corrupt downstream reasoning.
3.2 Perception Graph Model
Figure 3: Perception Graph generation: Given a video frame and its contextual description, generate the corresponding perception graph at time step t.
The perception graph model is a symbolic representation of perception, simulating how humans perceive, interpret, and prioritize information. It collects, structures, and stores scene data that enables formal reasoning later. The graph leverages pre-trained VLMs to approximate human interpretation of sensory input. Given AR imagery and task context, the model fuses multimodal signals into a compact graph and embeds their semantics in a latent perceptual space.
A perception graph consists of nodes and edges. Objects are represented as nodes, and object relationships are represented as edges. Given the context information and prompt (shown in the Figure  3), the VLM would extract the related objects with their relationships in the scene. At timestep $t$, the Perception Graph $G_{t}$ is generated as follows.
Graph Construction.
At discrete timestep $t\!\in\!\{1,\dots,T\}$ we ingest
$I_{t}\in\mathbb{R}^{H\times W\times 3},\qquad c_{t}\in\mathcal{C},$ (1)
where $I_{t}$ is the AR video frame and $c_{t}$ encodes task context. A pre-trained vision-language module $\mathcal{F}_{\text{VLM}}$ lifts the frame into a spatio-temporal perception graph
$G_{t}=(\mathcal{V}_{t},\mathcal{E}_{t})=\mathcal{F}_{\text{VLM}}\!\bigl{(}I_{t},\,c_{t}\bigr{)},$ (2)
with nodes $\mathcal{V}_{t}$ and edges $\mathcal{E}_{t}$:
$\mathcal{V}_{t}=\{v_{i}^{t}\}_{i=1}^{N_{t}},\qquad\mathcal{E}_{t}=\{\,e_{ij}^{t}\mid v_{i}^{t},v_{j}^{t}\in\mathcal{V}_{t}\}.$ (3)
For reasoning that needs the entire observation history (e.g. long-range causal inference) we define GraphSet
$\hat{G}_{t}=(\mathcal{V}_{1:t},\,\mathcal{E}_{1:t}),\quad\mathcal{V}_{1:t}=\bigcup_{s=1}^{t}\mathcal{V}_{s},\quad\mathcal{E}_{1:t}=\bigcup_{s=1}^{t}\mathcal{E}_{s}.$ (4)
Thus the GraphSet $\hat{G}_{t}$ is the union of all spatial and temporal connections that stitch object instances across time steps 1:t1{:}t. Each node still carries its original timestamp to preserve the temporal ordering.
Each node in the perception graph is represented by several attributes, including its ID, attack type, and timestep, along with a semantic description, visual features such as segmentation masks, a recognition probability score, contextual importance based on task relevance, and a reasonability score ranging from 3-3 to +3+3. These attributes are encoded using pre-trained vision-language and segmentation models to capture multimodal object representations. Each edge connects two nodes and includes similar parameters (node IDs, attack type, and timestep), along with a textual description capturing semantic, functional, navigational, or instructional relationships. Edge importance is derived from the contextual salience of the connected nodes.
NodeSet and EdgeSet.
After defining the temporal graph structure $\hat{G}_{t}$, which links perception graphs across time, we introduce the NodeSet and EdgeSet to represent the temporally connected nodes and edges, respectively.
At each time step $t$, the NodeSet for object $i$ is denoted as $\mathcal{N}_{i,t}$, and the collection of all such NodeSets is represented as the NodeSet List $\mathcal{N}_{t}$.
$\mathcal{N}_{i,t}$ $=\{\,v_{i}^{k}\mid 1\leq k\leq t\,\},\qquad i\in\mathcal{I}_{t},$ (5)
$\mathcal{N}_{t}$ $=\{\,\mathcal{N}_{i,t}\mid i\in\mathcal{I}_{t}\,\}.$
where $\mathcal{I}_{t}=\bigcup_{s=1}^{t}\{\,i\mid v_{i}^{s}\in\mathcal{V}_{s}\}$ is the set of all object IDs observed so far.
Similarly, EdgeSet and EdgeSet List is defined as:
$E_{ij,t}$ $=\{\,e_{ij}^{k}\mid 1\leq k\leq t\,\},\qquad i,j\in\mathcal{I}_{t},$ (6)
$E_{t}$ $=\{\,E_{ij,t}\mid i,j\in\mathcal{I}_{t}\,\}.$
Thus the GraphSet $\hat{G}_{t}$ can also represented by NodeSet and EdgeSet:
$\hat{G}_{t}=\bigl{(}\mathcal{V}_{1:t},\,\mathcal{E}_{1:t}\bigr{)},\quad\mathcal{V}_{1:t}=\!\!\bigcup_{i\in\mathcal{I}_{t}}\!\!\mathcal{N}_{i,t},\quad E_{1:t}=\!\!\bigcup_{i,j\in\mathcal{I}_{t}}\!\!\mathcal{E}_{ij,t}.$ (7)
3.3 Particle Filter Model
Our particle filter model tracks graph evolution and analyzes adversarial perturbations. It proceeds in three stages: prediction and matching, attack detection, and state estimation (cf. Fig. 4). Each entity (node or edge) is treated as a particle At each timestep, particles match the incoming observations with prior particles. The attack detection algorithm applies statistical tests to newly matched particles and discards those marked as attacked. Finally, weighting, resampling, and estimation produces an updated, temporally consistent graph state, preserving persistent NodeSet/EdgeSet identities and the relational structure over time.
Figure 4: The particle filter-based attack detection framework consists of three main modules: matching, attack detection, and state estimation.
Prediction and Matching.
To maintain persistent object identities across frames, we equip the NodeSet with the Prediction-Matching algorithm and use it to decide whether a newly detected Node belongs to an existing object list or creates a new one.
For any node $v_{i}^{t}$ we denote $\mathbf{n}(v)$ = name embedding, $\mathbf{d}(v)$ = description embedding, $\mathbf{f}(v)$ = visual feature embedding, all $\in\mathbb{R}^{d_{m}}$. Cosine similarity for modality $m\in\{n,d,f\}$ is defined as $\operatorname{sim}_{m}(u,v)$.
Let $\mathcal{A}_{t-1}$ be the set of NodeSet IDs that are still active at time t1t-1. Each set maintains an Active Level $A_{i}^{t}\in[0,3]$ (initialized to $3$) indicating recent presence: if unmatched at time $t$, $A_{i}^{t}\leftarrow\max(0,A_{i}^{t-1}-1)$; if matched, $A_{i}^{t}\leftarrow\min(3,A_{i}^{t-1}+1)$. When $A_{i}^{t}=0$, the node is considered absent and ceases matching. For every active object ID $i\in\mathcal{A}_{t-1}$ and each modality $m\in\{n,d,f\}$ compute the mean and standard deviation of pairwise similarities inside the NodeSet:
$\mu_{i,m}^{\,t-1}$ $=\frac{2}{|\mathcal{N}_{i,{t-1}}|(|\mathcal{N}_{i,{t-1}}|-1)}\sum_{\begin{subarray}{c}p<q\\ v_{i}^{p},v_{i}^{q}\in\mathcal{N}_{i,{t-1}}\end{subarray}}\operatorname{sim}_{m}(v_{i}^{p},v_{i}^{q}),$ (8)
$\sigma_{i,m}^{\,t-1}$ $=\mathrm{StdDev}\!\left\{\operatorname{sim}_{m}(v_{i}^{p},v_{i}^{q})\;\big{|}\;p<q,\;v_{i}^{p},v_{i}^{q}\in\mathcal{N}_{i,t-1}\right\}.$
To match the observed node to NodeSet, given a fresh node $v_{j}^{t}$ and a candidate NodeSet $i$, define
$s_{j,i,m}\;=\;\frac{1}{|\mathcal{N}_{i,t-1}|}\sum_{v\in\mathcal{N}_{i,t-1}}\operatorname{sim}_{m}(v_{j}^{t},v),\qquad m\in\{n,d,f\}.$ (9)
Candidate filter ($\leq 1$std in every modality):
$\mathcal{C}_{j}\;=\;\Bigl{\{}i\in\mathcal{A}_{t-1}\;\Bigl{|}\;\bigl{|}s_{j,i,m}-\mu_{i,m}^{\,t-1}\bigr{|}\leq\sigma_{i,m}^{\,t-1},\;\forall m\in\{n,d,f\}\Bigr{\}}.$ (10)
For each remaining candidate, measure the distance
$D_{j,i}\;=\;\sum_{m\in\{n,d,f\}}\bigl{|}s_{j,i,m}-\mu_{i,m}^{\,t-1}\bigr{|}.$ (11)
Finally, the assignment rule:
$\mu_{t}(v_{j,t})=\begin{cases}\displaystyle\arg\min_{i\in\mathcal{C}_{j}}D_{j,i},&\mathcal{C}_{j}\neq\varnothing,\\[6.0pt] \text{new\_id},&\mathcal{C}_{j}=\varnothing.\end{cases}$ (12)
If $\mu_{t}(v_{j}^{t})$ returns an existing ID, append $v_{j}^{t}$ to that NodeSet and update $A_{i}^{t}$; otherwise spawn a new NodeSet. Confirmed edges $e_{ab}^{t}$ are then inserted into $E_{\mu_{t}(a)\,\mu_{t}(b),t}$ as before.
After matching, the NodeSet goes into the Attack detection process.
Attack Detection.
To detect text (Type 1) and visual (Type 2) modification attacks, those that manifest as anomalies in embedding space, we employ a dedicated statistical procedure. Because the same description embeddings are later reused when assessing EdgeSets, we illustrate the method with the NodeSet case; the extension to edges is entirely analogous. A NodeSet with a fresh match in frame $t$ is denoted $\mathcal{N}_{i,t}=\mathcal{N}_{i,t-1}\cup\{v_{i}^{t}\}$. Let $\tilde{\mathcal{N}}_{i,t-1}\subseteq\mathcal{N}_{i,t-1}$ be the subset of nodes that have not been flagged as attacks in earlier frames; these constitute the “clean” reference set.
For each modality $m\!\in\!\{d,f\}$ (description embedding, visual feature embedding), we use the Principal Component Analysis (PCA) to project it to a low-dimensional space, which removes noisy, low-variance directions. A PCA mapping on the clean reference data is computed by
$\mathbf{W}_{i,m}^{\,t-1}\;=\;\operatorname{PCA}\!\bigl{(}\{\mathbf{m}(v)\}_{v\in\tilde{\mathcal{N}}_{i,t-1}},\;k_{m}\bigr{)},$ (13)
where $k_{m}$ is chosen such that $\geq\!99\%$ variance is retained. The projected embeddings are
$\mathbf{z}_{m}(v)=\mathbf{W}_{i,m}^{\,t-1\,\top}\mathbf{m}(v)\;\in\mathbb{R}^{k_{m}}.$ (14)
For each modality, fit the Gaussian reference model
$\mathbf{z}_{m}(v)\;\sim\;\mathcal{N}\!\bigl{(}\mathbf{c}_{i,m}^{\,t-1},\,\boldsymbol{\Sigma}_{i,m}^{\,t-1}\bigr{)},\quad v\in\tilde{\mathcal{N}}_{i,t-1},$ (15)
with the centroid and covariance of PCA-reduced embeddings
$\mathbf{c}_{i,m}^{\,t-1}=\frac{1}{|\tilde{\mathcal{N}}_{i,t-1}|}\sum_{v\in\tilde{\mathcal{N}}_{i,t-1}}\mathbf{z}_{m}(v),$ (16)
$\boldsymbol{\Sigma}_{i,m}^{\,t-1}=\mathrm{Cov}\!\bigl{\{}\mathbf{z}_{m}(v)\bigr{\}}.$ (17)
Use the Mahalanobis distance for measuring Text and Visual Modification attack. For a new node $v_{i}^{t}$ compute
$\mathcal{M}_{i,t}^{(m)}\;=\;\bigl{(}\mathbf{z}_{m}(v_{i}^{t})-\mathbf{c}_{i,m}^{\,t-1}\bigr{)}^{\!\top}\bigl{(}\boldsymbol{\Sigma}_{i,m}^{\,t-1}\bigr{)}^{-1}\bigl{(}\mathbf{z}_{m}(v_{i}^{t})-\mathbf{c}_{i,m}^{\,t-1}\bigr{)},$ (18)
Assume under the no-attack hypothesis, the squared Mahalanobis statistic $\mathcal{M}_{i,t}^{(\cdot)}\sim\chi^{2}_{k,\alpha}$, which is the $(1-\alpha)$-quantile of a chi-square distribution with $k$ degrees of freedom. So flag the attack as
$v_{i}^{t}\;\text{as}\;\begin{cases}\text{Type 1 Attack(text)}&\text{if }\mathcal{M}_{i,t}^{(d)}>\chi^{2}_{k_{d},\alpha},\\[2.0pt] \text{Type 2 Attack(image)}&\text{if }\mathcal{M}_{i,t}^{(f)}>\chi^{2}_{k_{f},\alpha}.\end{cases}$ (19)
For the structural attacks, specifically object removal (Type 3) and object addition (Type 4), the detector relies on lightweight temporal heuristics rather than statistical distance. After each matching step the algorithm keeps a two-frame sliding window of NodeSet activity. If an object that previously carried a high task-importance score ($\pi\geq\pi_{\text{high}}$) fails to obtain a match in both the current and the immediately preceding frame, the system labels the disappearance a removal attack. Conversely, any NodeSet that first appears within the last two frames and whose reasonability score falls below a predefined threshold ($\rho\leq\rho_{\text{low}}$) is treated as an addition attack.
In the attack detection final step, if $v_{i}^{t}$ is flagged as attack, it is appended to the $\mathcal{N}_{i,t}$ but excluded from future Gaussian updates; otherwise going to the Node Estimation process, including weight and resample, then stored in a clean reference $\tilde{\mathcal{N}}_{i,t}$.
Estimation.
After a new observation $v_{i,t}$ has been matched to object $i$ and detected as a non-attack node, the particle filter estimates the new node by refining that observation by resampling historical detections that resemble it. The estimation proceeds as follows.
Let the newly matched observation of NodeSet $i$ at frame $t$ be $v_{i,t}$ and denote by
$\tilde{\mathcal{N}}_{i,t-1}=\{v^{(1)},\dots,v^{(M)}\}\subseteq\mathcal{N}_{i,t-1},\quad M\geq 5,$
the set of non-attack detections accumulated up to t1t\!-\!1. Considering each attribute as a particle, the filter updates $v_{i,t}$ by weighted resampling over this reference set.
For any scalar attribute $a\in\{p,\pi,\rho\}$ (probability, importance, reasonability), let $a^{\star}=a(v_{i}^{t})$ and $a^{(j)}=a(v^{(j)})$ ($j=1\dots M$). Define the weight
$\sigma_{a}=\max\Bigl{(}\operatorname{Std}\{a^{(1:M)}\},\;\sigma_{\min}\Bigr{)},\qquad$ (20)
$w_{j}=\frac{\exp\!\bigl{[}-(a^{(j)}-a^{\star})^{2}/(2\sigma_{a}^{2})\bigr{]}}{\sum_{k=1}^{M}\exp\!\bigl{[}-(a^{(k)}-a^{\star})^{2}/(2\sigma_{a}^{2})\bigr{]}}.$ (21)
So $w_{j}$ is the normalized weight of the particle, based on its distance to the new particle. Then use the weight to resample $M$ indices $s_{1},\dots,s_{M}\sim\operatorname{Multinomial}(w_{1:M})$ yields
$\hat{a}\;=\;\frac{1}{M}\sum_{k=1}^{M}a^{(s_{k})},$ (22)
where $\hat{a}$ is the estimated new scalar particle.
For embedding attributes, modality $m\in\{n,d,f\}$ (name, description, image) let $\mathbf{m}^{\star}=\mathbf{m}(v_{i}^{t})$ and $\mathbf{m}^{(j)}=\mathbf{m}(v^{(j)})$. Cosine similarity gives
$sim_{j}=\frac{\mathbf{m}^{(j)}\!\cdot\!\mathbf{m}^{\star}}{\|\mathbf{m}^{(j)}\|\,\|\mathbf{m}^{\star}\|},\qquad w_{j}=\frac{\exp(\lambda_{m}sim_{j})}{\sum_{k=1}^{M}\exp(\lambda_{m}sim_{k})},$ (23)
with a scale factor $\lambda_{m}>0$. Resampling as above yields indices $s_{1},\dots,s_{M}\sim\operatorname{Multinomial}(w_{1:M})$ and the unit-normalised mean
$\hat{\mathbf{m}}\;=\;\frac{\sum_{k=1}^{M}\mathbf{m}^{(s_{k})}}{\Bigl{\|}\sum_{k=1}^{M}\mathbf{m}^{(s_{k})}\Bigr{\|}},$ (24)
would be the resampled embedding particle.
Finally, replacing each raw attribute in $v_{i}^{t}$ by its resampled estimate gives the estimated node
$\hat{v}_{i,t}=\bigl{(}\hat{p},\;\hat{\pi},\;\hat{\rho},\;\hat{\mathbf{n}},\;\hat{\mathbf{d}},\;\hat{\mathbf{f}}\bigr{)}$, (25)
which becomes the particle filter’s estimation state for object $i$ at time $t$. If $M<5$ (insufficient history) we skip resampling and simply set $\hat{v}_{i,t}=v_{i,t}$.
4 Experiment
Table 1: Results of the state-of-the-art methods and CADAR on Extended AR-VIM Dataset
Method Attack Detection: Accuracy/F1-score (%) Localization Acc. of attack(%)
Text Visual Removal Addition Text+Visual Add.+Rem. Overall
CADAR (Ours) 80.2/ 74.9 82.3/ 76.1 83.0/ 86.3 82.1/ 83.6 67.0/ 69.7 86.1/ 75.2 80.1/77.6 68.0
GPT-o3 62.5/56.6 23.8/29.4 76.2/66.7 61.9/55.3 16.7/25.0 14.3/25.0 39.2/39.8 -
Gemini 2.5 66.7/71.8 66.7/77.8 81.0/69.4 90.5/88.4 14.3/18.1 71.4/71.4 69.4/71.1 -
3D ResNet 12.5/22.2 57.1/72.7 57.1/34.7 42.8/31.6 - - 42.4/40.3 -
ViViT 50.0/44.4 28.6/33.3 71.4/47.6 14.3/20.0 - - 41.1/36.3 -
In this section, we provide a detailed description of the CADAR implementation and the experimental settings used for our evaluation. We then present and analyze the numerical results of our experiments, demonstrating that CADAR significantly outperforms several baselines, achieving an overall accuracy of 80.1% compared to 69.4% for the next-best model. An ablation study is also conducted to analyze the contributions of our key components.
4.1 Implementation of CADAR
CADAR is implemented using a modular architecture. The perception graph is constructed and managed through several key components. The Node.py and Edge.py modules generate multimodal object representations by combining textual and visual features: GPT-4.1 is used for generating descriptive text, google/owlv2-base-patch16-ensemble (Minderer, Gritsenko, and Houlsby 2024) enables zero-shot object detection, and SAM2 (Ravi et al. 2024) provides object segmentation. Embeddings are extracted using all-MiniLM-L6-v2 for object names, all-mpnet-base-v2 for descriptions, and ResNet-34 for image features. The Graph.py module manages the perception graph structure, maintaining node–edge relationships and temporal information. Temporal reasoning is handled by NodeSet.py and EdgeSet.py, which update the graph across frames, detect potential attacks, and evolve the graph state. Attack detection is based on statistical anomaly detection using PCA to preserve 99% of variance, with thresholds computed from a 99% confidence level of the chi-squared distribution. Finally, Utils.py provides supporting tools for video processing and mathematical computations.
4.2 Experimental Settings
Extended AR-VIM Dataset.
We base our experiments on an extended version of the AR-VIM dataset, which is specifically designed to evaluate visual information manipulation in AR environments (Xiu and Gorlatova 2025). The Extended AR-VIM is the first publicly available AR cognitive attack dataset (details in Appendix). Each sample in the dataset consists of a short video collected by an AR device (Meta Quest 3 and Samsung Galaxy S24), approximately 5-10 seconds in length, where virtual content appears at a specific frame to trigger an attack. The dataset has 160 videos in different scenes, spanning all four cognitive-attack classes. The scene covers transportation (e.g., highway, train, aircraft), experimental settings (e.g., lab equipment), industrial environments (e.g., construction, agriculture), workspaces (e.g., warehouse, office), as well as medical, daily-life, and public-safety domains. These range from altering freeway exit-ramp signage to suppressing voltage warnings on live equipment. Together, these videos provide a diverse, realistic, and tightly controlled test bed for evaluating the perception-level of cognitive attacks in AR environments. The extended AR-VIM dataset consists of 41 removal attack samples, 35 addition attack samples, 40 text manipulation attack samples, 30 visual manipulation attack samples, and 14 mix-class attack (text + visual manipulation attacks and addition + removal attacks) samples. For baseline training models, it also contains 40 non-attacked videos.
Considering the scene availability, the dataset was collected using two complementary methods: (1) The monitor-based method places AR content on a monitor screen with background images and captures it via video recording, allowing rapid and controlled generation of attack instances; (2) The real-world-based method, in contrast, uses AR headsets to capture first-person views in physical environments, introducing natural visual noise and depth information. Employing both methods is critical to capturing the trade-off between efficient data generation and real-world fidelity.
Task and Metrics.
We evaluate our model on the task of attack classification in videos. CADAR supports frame-wise detection, variable-length multi-label outputs, and localization of multiple simultaneous attacks, which are capabilities that few existing models possess. As most baselines can only classify attacks at the video level and lack the capability for fine-grained, frame-level detection or localization, we compare CADAR against both advanced reasoning VLMs (GPT-o3 and Google Gemini 2.5 (Comanici et al. 2025)) and trained video classification models, including 3D-ResNet (Tran et al. 2018) and ViViT (Arnab et al. 2021). We use accuracy and F1-score to evaluate attack detection performance: frame-wise for CADAR, and video-level for the baselines. Additionally, we report attack localization accuracy, which measures how precisely the model locates the attacks in the frame. Results are summarized in Table 1. For mixed-attack instances, a detection is considered correct only if both attack types are correctly identified.
4.3 Numerical Results and discussion
Table 1 presents the evaluation results for attack classification on the extended AR-VIM dataset. CADAR outperforms both the supervised learning baselines and the pre-trained VLMs. Among the different attack types, removal and addition attacks are generally easier to detect, leading to higher accuracy compared to text and visual modifications. Specifically, CADAR achieves a detection accuracy of 83.0% for removal attacks and 82.1% for addition attacks, compared to 80.2% for text and 82.3% for visual attacks. Notably, CADAR demonstrates balanced and robust performance across all attack categories, achieving superior accuracy and F1 scores. For instance, its overall accuracy/F1 score is 80.1%/77.6%, significantly higher than the next best model, Gemini 2.5, which scores 69.4%/71.1%.
Supervised models such as 3D-ResNet and ViViT require large-scale training data and perform poorly in zero- or few-shot scenarios. For example, 3D-ResNet achieves an overall accuracy of 42.4%, and ViViT achieves 41.1%. In contrast, pre-trained VLMs leverage prior knowledge and exhibit strong reasoning abilities, but their outputs are often unstable and prone to hallucination. For example, GPT-o3 and Gemini 2.5 achieve overall accuracies of 39.2% and 69.4%, respectively.
CADAR offers a novel approach that addresses these limitations and achieves the best overall performance. However, it is not without shortcomings. The localization accuracy is constrained by the capabilities of the zero-shot object detection and segmentation models, which struggle in complex scenes or with unfamiliar objects. CADAR achieves a localization accuracy of 68.0%. Moreover, reduced performance on text attacks suggests that highly subtle edits, e.g., a single-character change that scarcely alters semantics or appearance, remain difficult for the detector.
4.4 Ablation Study
Estimation Module.
Figure 5: Ablation study on estimation module and reference frames.
Fig. 5(a) shows our ablation study examining the impact of the estimation module in particle filtering. When a polluted node that might have bad data due to VLM unstable output or due to an undetected attack, the estimation module’s weighting and resampling can effectively prevent it from corrupting the ground truth knowledge. In the ablation study, we choose 12 videos for all attack types. As shown in the left plot of the figure, for text and visual modification attacks, the detection accuracy drops significantly as the number of frames without the estimation module increases. Specifically, the accuracy for text attacks decreases from approximately $83\%$ to $69\%$ and from approximately $81\%$ to $55\%$ for visual attacks. This indicates that the estimation module is crucial for correctly identifying and mitigating errors in more subtle attack types. In contrast, for removal and addition attacks, the accuracy remains relatively stable, suggesting these attacks are more straightforward to detect and less dependent on the estimation module’s corrective capabilities.
Reference Frames.
Fig. 5(b) shows the impact of the number of reference frames on detection accuracy. A larger number of reference frames, or more prior knowledge of the scene, will result in higher accuracy. For text and visual modification attacks, increasing the number of reference frames from 3 to 11 significantly improves performance. The accuracy for text attacks increases from around $55\%$ to $80\%$ and from around $50\%$ to $78\%$ for visual attacks. This demonstrates that a larger set of reference frames provides more prior knowledge for more accurate statistical reasoning, which is essential for detecting subtle text and visual changes. Similar to the estimation module study, the detection accuracy for removal and addition attacks remains largely unaffected by the number of reference frames, consistently staying above $80\%$. This reinforces the finding that these attacks are less sensitive to the amount of contextual information and prior knowledge.
5 Conclusion
We presented CADAR, a neuro-symbolic framework that couples a perception graph with particle-filter reasoning to yield interpretable attack detection in AR/MR. On the Extended AR-VIM dataset, CADAR performs strongly across all attack types, and ablations highlight the value of our estimation module and richer reference histories. Limitations stem from the underlying open-vocabulary detection/segmentation, and miss very fine-grained augmentations. Future work will pursue task-specific fine-tuning of perception models, and larger-scale real-world evaluations and user studies. Overall, CADAR highlights the value of neuro-symbolic design by combining the adaptability of foundation models with symbolic reasoning for a robust and interpretable AR safety mechanism.
Appendix A Extended AR-VIM Dataset Samples
Representative sample videos for each attack class (excluding the No Attack category) is shown below. The full dataset consists of 160 videos across diverse scenes.
Figure 6: Sample dataset content for each AR Attack Class.
Appendix B Dataset Construction
For the real-world scenes in the Extended AR-VIM dataset, we collected additional samples using a Meta Quest 3 headset. The data collection setup is shown below
Figure 7: (a) First-person view from the AR headset: the virtual object (e.g. a dog and textual cues) is interactive and can be toggled on or off using the “Button B” on the right controller, while “Button A” starts or stops video recording. (b) Third-person view showing the data collector wearing the Meta Quest 3 headset and interacting with the AR environment using handheld controllers.
Appendix C Model Architectures and Training
For our supervised-learning baselines—ResNet and ViViT. Training hyper-parameters are shown below. We conducted all training and evaluation using Google Colab with NVIDIA A100 GPUs (40GB memory). The experiments were run in a Linux environment via Colab’s remote infrastructure, conducted using PyTorch 2.6.0 with Python 3.11.
3D ResNet ViViT
Value Range # Tried Value Range # Tried
initial/end lr 5e-5 [5e-5, 5e-3] 5 1e-4 [5e-5, 5e-3] 4
weight decay 5e-3 [1e-4, 1e-1] 5 5e-3 [1e-3, 1e-1] 4
batch size 20 [8–20] 3 16 [4–16] 2
epochs 10 [5–50] 6 10 [5–50] 4
scheduler StepLR Cosine
optimizer AdamW AdamW
Table 2: Training hyper-parameters, tested ranges, and number of values tried for ViViT and 3D ResNet.
We use ViViT-B/16x2 (pretrained on Kinetics-400) as the backbone for our ViViT model, and a 3D ResNet-18 for the ResNet baseline. Both models are trained on the full training set using inverse-frequency sampling to balance class distributions. The classification task involves five classes: the four attack types, along with an additional No Attack category. All input videos are resized to 400×400 pixels and normalized using a mean of [0.45, 0.45, 0.45] and a standard deviation of [0.225, 0.225, 0.225]. Each input consists of eight uniformly sampled frames from a 13-second clip. We use Focal Loss with a gamma of 2.0 as the training objective. Mixed precision (FP16) training is enabled via HuggingFace Accelerate for both models. ViViT is trained with a cosine learning rate scheduler, while the 3D ResNet uses a StepLR scheduler (step size = 10, gamma = 0.1). Final hyperparameter configurations for both models were selected based on the settings that yielded the highest validation accuracy.
Appendix D Benchmark Prompt Templates (GPT-o3 and Gemini 2.5)
For our baseline VLM models, we use the following prompt for both GPT-o3 and Google Gemini 2.5. Related work include(Zhang et al. 2025; Zou et al. 2024; Qiao et al. 2024; Zhang et al. 2024b; Jing et al. 2024; Gao et al. 2024; Ravari et al. 2024; Zhang, Aggarwal, and Lan 2025; Yu et al. ; Zhang, Imani, and Lan 2024)
You are an AR video attack analyst. Your job is to identify the presence of 5 different kinds of AR attacks in the video. The definitions for each attack are as seen below: A text modification attack manipulates textual information embedded in the environment, thereby changing the semantic message conveyed by the scene. For example, replacing a NO PARKING FIRE LANE sign with one reading FREE PARKING FIRE LANE, for instance, directly reverses the intended directive and can create significant safety hazards. Such attacks are especially consequential in contexts where text offers primary contextual or navigational cues.
A visual modification attack alters an object’s appearance—its color, shape, or iconography—or repositions it in an implausible manner. Appearance changes, such as transforming a pedestrian-crossing sign into a bicycle-crossing sign or switching a green traffic signal to red, degrade recognition accuracy and mislead both human users and perception systems. Spatial changes include shifting an object away from its logical location or removing it entirely; for example, relocating a stop sign to the centre of an intersection. These manipulations can be quantified by analysing pixel-level discrepancies and deviations from expected spatial distributions.
A removal attack hides or deletes critical objects, thereby eliminating essential information and distorting scene semantics. In a fire evacuation scenario, a blank patch that obscures an emergency-exit sign on a door. By erasing this safety-critical cue, the attack deprives occupants of clear evacuation guidance, potentially endangering lives during an emergency. By analyzing nodes from the scene graph, removal attacks disrupt relational structure and semantic integrity.
An addition attack introduces new, contextually inconsistent, or misleading objects into the scene, thereby distorting its intended semantics or altering user perception. For example, a fake road sign indicating a detour is inserted into a clear street, potentially redirecting traffic and causing confusion. These attacks may include adding fictitious warning signs, virtual obstructions, or irrelevant icons that do not exist in the physical environment. Such manipulations exploit the overlay nature of AR systems and are particularly dangerous when added elements resemble legitimate cues, misleading both human users and perception models. Detection relies on identifying unexpected objects that conflict with the scene’s established structure and context.
The following input will contain a video. Please output the attack kind. Note that there may be more than 1 kind of attack present in the video, or a video could have no attack present at all. Along with the attack, mention the object the attack was performed on or describe the attack briefly in the context of the video.

### v3

@@
 Abstract
-Augmented Reality (AR) enriches perception by overlaying virtual elements on the physical world.
-Due to its growing popularity, cognitive attacks that alter AR content to manipulate users’ semantic perception have received increasing attention.
-Existing detection methods often focus on visual changes, which are restricted to pixel- or image-level processing and lack semantic reasoning capabilities, or they rely on pre-trained vision–language models (VLMs), which function as black-box approaches with limited interpretability.
-In this paper, we present CADAR, a novel neurosymbolic approach for cognitive attack detection in AR.
-It fuses multimodal vision–language inputs using neural VLMs to obtain a symbolic perception-graph representation, incorporating prior knowledge, salience weighting, and temporal correlations.
-The model then enables particle-filter based statistical reasoning–a sequential Monte Carlo method–to detect cognitive attacks.
-Thus, CADAR inherits the adaptability of pre-trained VLM and the interpretability and reasoning rigor of particle filtering.
-Experiments on an extended AR cognitive attack dataset show accuracy improvements of up to $10.7\%$ over strong baselines on challenging AR attack scenarios, underscoring the promise of neurosymbolic methods for effective and interpretable cognitive attack detection.
+Augmented Reality (AR) enriches human perception by overlaying virtual elements onto the physical world.
+However, this tight coupling between virtual and real content makes AR vulnerable to cognitive attacks: manipulations that distort users’ semantic understanding of the environment.
+Existing detection methods largely focus on visual inconsistencies at the pixel or image level, offering limited semantic reasoning or interpretability.
+To address these limitations, we introduce CADAR, a neuro-symbolic framework for cognitive attack detection in AR that integrates neural and symbolic reasoning.
+CADAR fuses multimodal vision–language representations from pre-trained models into a perception graph that captures objects, relations, and temporal contextual salience.
+Building on this structure, a particle-filter-based statistical reasoning module infers anomalies in semantic dynamics to reveal cognitive attacks.
+This combination provides both the adaptability of modern vision–language models and the interpretability of probabilistic symbolic reasoning.
+Preliminary experiments on an AR cognitive-attack dataset demonstrate consistent advantages over existing approaches, highlighting the potential of neuro-symbolic methods for robust and interpretable AR security.
 1 Introduction
-Augmented Reality (AR) and Mixed Reality (MR) technologies have rapidly advanced in recent years, offering users an immersive perceptual experience by overlaying contextually relevant information and virtual objects onto their physical environment (Zheng and Yuan 2023).
-As AR/MR are rapidly changing how users perceive, interpret, and interact with the world around them, new vulnerabilities have emerged and attracted increasing research attention.
-In particular, the reliance on visual and spatial cues in AR/MR creates an implicit trust in the accuracy and authenticity of overlaid content.
-This trust can be exploited by adversaries to launch cognitive attacks, which use malicious manipulations of virtual content to mislead, distract, or coerce users without their awareness (Xiu, Scargill, and Gorlatova 2025; Cheng et al.
-2023; Teymourian et al.
-2025).
-Unlike traditional cybersecurity threats that target data or system integrity, cognitive attacks in AR/MR (such as the examples shown in Fig.
-1) operate at the perceptual level, manipulating users’ semantic interpretation through subtle or deceptive visual alterations (Xiu and Gorlatova 2025).
+Augmented Reality (AR) and Mixed Reality (MR) technologies increasingly overlay virtual content onto the physical world, reshaping how users perceive and interact with their surroundings [zheng2023review].
+This tight integration creates new vulnerabilities, as users tend to trust overlaid visual and spatial cues.
+Adversaries can exploit this trust to launch cognitive attacks, where malicious virtual content is injected or altered to mislead, distract, or manipulate users without their awareness [10924575, cheng2023exploring, teymourian2025sok, xiu2025detectingvisualinformationmanipulation].
+Unlike traditional cybersecurity threats that target system integrity, cognitive attacks operate at the perceptual level, subtly manipulating users’ semantic understanding of a scene (examples in Fig.1).
 Figure 1: Illustrative AR/MR cognitive attacks.
@@
 (b) Visual Modification attack.
-(c) Removal Attack.
-(d) Addition Attack.
-Prior work on detecting such cognitive attacks often employs computer vision techniques to identify visual changes or adversarial distortions (Qiu et al.
-2019; Duan et al.
-2022a, b; Li et al.
-2024; Acharya et al.
-2022).
+(c) Obstruction Attack.
+(d) Injection Attack.
+Prior work on detecting anomaly and attacks often employs computer vision techniques to identify visual changes or adversarial distortions [qiu2019review, gu2024anomalygpt, qu2023towards, wang2022objectformer, duan2022confusing, li2024survey, acharya2022detecting].
 However, these methods rely on pixel/image processing, lacking the ability to model the human perception process that not only involves prior knowledge and contextual information, but also selectively focuses on aspects of the visual field for semantical interpretation.
-As a result, these methods may struggle to capture smart cognitive attacks manipulating only a small number of pixels and also could suffer from false positives due to benign visual perturbations that do not affect users’ semantic interpretation (Li et al.
-2024).
-More recent studies propose pre-trained vision–language models (VLMs) (Xiu, Scargill, and Gorlatova 2025; Roy et al.
-2025) to achieve semantic-level detection, by processing visual features and incorporating human-like knowledge.
+As a result, these methods may struggle to capture smart cognitive attacks manipulating only a small number of pixels and also may suffer from false positives due to benign visual perturbations that do not affect users’ semantic interpretation [li2024survey].
+More recent studies propose pre-trained vision–language models (VLMs) [xiu2025viddar, roy2025zero] can achieve semantic-level detection by processing visual features and incorporating human-like knowledge.
 But the reasoning of pre-trained VLMs is a black-box approach offering limited interpretability.
-It is hard to leverage domain-specific knowledge or model scene/context evolutions or temporal correlations in cognitive attack detection.
+It is difficult to leverage domain-specific knowledge or model scene/context evolutions or temporal correlations in cognitive attack detection.
 In this paper, we consider cognitive attacks that aim to manipulate users’ semantic perception in AR/MR.
-We propose CADAR (Cognitive Attack Detection in Augmented Reality), a neuro-symbolic approach encompassing two key components working in concert: (i) a symbolic perception-graph model, which is extracted using pre-trained neural VLM models to represent scene/context evolution at semantic-level, ; and (ii) a particle-filter model, which is commonly used in non-linear and stochastic systems.
+We propose CADAR (Cognitive Attack Detection in Augmented Reality), a neuro-symbolic approach encompassing two key components: (i) a symbolic perception-graph model, inspired by knowledge and scene graphs [hogan2021knowledge, chang2021comprehensive, yun2025missiongnn, shah2022lmnavrobotic].
+It extracted using pre-trained neural VLM models to represent scene/context evolution at the semantic-level, and (ii) a particle-filter model, which is commonly used in non-linear and stochastic systems.
 It applies a sequential Monte Carlo method to the perception graph for reasoning and detecting cognitive attacks, stabilizes inherently noisy or inconsistent VLM outputs, and mitigates errors or missed detections through probabilistic fusion.
-It also offers granular, flexible temporal detection: by using reference sequences, frame-wise inference localizes attacks per frame, to produce a variable number of labels when needed.
-Thus, the CADAR approach inherits the adaptability of pre-trained VLM (neural part), as well as the interpretability and reasoning rigor of particle filtering (symbolic part).
-It semantically interprets and models the sequential scene changes for robust cognitive attacks detection in AR/MR, overcoming limitations in existing work.
-Our key contributions are as follows:
-  • Neuro-symbolic modeling and reasoning.
-We present a neuro-symbolic framework tailored to AR/MR cognitive attack detection.
-A pre-trained neural VLM generates symbolic perception graphs that stores objects, attributes, and relations for explicit, interpretable reasoning.
-  • Perception graph representation.
-We propose a perception graph schema that formally encodes multi-modal scene/context evolution at the perceptual level, integrating outputs from complementary modules (image and text encoders, object detection, and segmentation) for cognitive attack inference.
-  • Particle-filter-based statistical reasoning.
-We are the first to apply particle filtering to symbolic perception graphs.
-Not only does our method improve the robustness of stochastic VLM outputs, mitigating noise and hallucination, but it also enables interpretable statistical procedures to reason about cognitive attacks.
-  • We create the Extended AR-VIM dataset, which is the first publicly available AR cognitive attack dataset, by expanding the existing AR-VIM dataset (Xiu and Gorlatova 2025).
-CADAR evaluation shows it significantly outperforms several baselines, achieving an overall accuracy of 80.1% compared to 69.4% for the next-best model.
-The rest of the paper is organized as follows: Section 2 reviews related work; Section 3 presents our method; Section 4 reports experimental results and Section 5 concludes the paper.
+It also offers granular, flexible temporal detection, using reference sequences, frame-wise inference localizes attacks per frame, to produce a variable number of labels when needed.
+Thus, the CADAR approach inherits the adaptability of the pre-trained VLM (neural part), as well as the interpretability and reasoning rigor of particle filtering (symbolic part).
+It semantically interprets and models the sequential scene changes for robust cognitive attack detection in AR/MR, overcoming existing work’s limitations.
+We summarize our contributions as follows:
+  • Neuro-symbolic framework. We introduce a hybrid architecture that converts VLM outputs into symbolic perception graphs for explicit and interpretable reasoning in AR cognitive attack detection.
+  • Perception graph schema. We design a spatio-temporal graph representation that integrates multimodal information from image/text encoders and detection modules to capture scene evolution.
+  • Particle-filter reasoning. We apply particle filtering to symbolic perception graphs, improving robustness to noisy VLM outputs and enabling principled statistical anomaly detection.
+  • CADAR-50K dataset. We build the first public AR cognitive attack dataset and demonstrate that CADAR achieves 74.6% accuracy, significantly outperforming all baseline methods.
+The rest of the paper is organized as follows: Section 2 reviews related work; Section 3 presents our method; Section 4 reports experimental results; and Section 5 concludes the paper.
 2 Related Work
 Cognitive Attacks in AR/MR.
-One well-studied category of visual attacks is adversarial attacks, which introduce imperceptible yet malicious perturbations to input data to mislead machine learning models (Zhang et al.
-2024a; Mumcu and Yilmaz 2024).
-In contrast, cognitive attacks target perception-level semantics and aim to alter a human user’s understanding of a scene (Cheng et al.
-2023; Xiu and Gorlatova 2025).
+One well-studied category of visual attacks is adversarial attacks, which introduce imperceptible yet malicious perturbations to input data to mislead machine learning models [zhang2024adversarial, mumcu2024multimodal].
+In contrast, cognitive attacks target perception-level semantics and aim to alter a human user’s understanding of a scene [cheng2023exploring, xiu2025detectingvisualinformationmanipulation].
 Augmented Reality (AR) systems, which overlay virtual content onto the physical world to enhance perception and decision-making, are particularly vulnerable to such attacks.
@@
 Existing methods often rely on access to the original, unaltered scene for comparison or employ pre-trained VLMs with limited interpretability.
-Knowledge Graphs.
-Graph-based representations offer a compact and expressive way to encode entities and the possibly complex relations among them, making graphs well-suited for downstream prediction and retrieval tasks.
-A knowledge graph represents each entity as a node and each semantic relation as an edge, thereby capturing structured world knowledge at scale (Hogan et al.
-2021).
-A scene graph extends this idea to visual understanding: it organizes objects, their attributes, and inter-object relationships into a graph that supports higher-level reasoning over images and videos (Chang et al.
-2021).
-Numerous task-specific graph formulations have also emerged, for example, mission-specific knowledge graphs for situational awareness (Yun et al.
-2025), topological graphs for robot navigation (Shah et al.
-2022), and interaction networks that model physical relations among objects (Battaglia et al.
-2016).
 Neuro-symbolic Methods.
-Neuro-symbolic approaches couple deep neural networks with symbolic reasoning–i.e., reasoning over explicitly represented knowledge in formal languages, to capitalize on the complementary strengths of subsymbolic pattern recognition and symbolic compositionality (Bhuyan et al.
-2024).
-Within knowledge-graph research, most neuro-symbolic work addresses link prediction or graph completion, where the hybrid models offer interpretable rules, guided learning, and more transparent decision pathways (DeLong, Mir, and Fleuriot 2024).
+Neuro-symbolic approaches couple deep neural networks with symbolic reasoning–i.e., reasoning over explicitly represented knowledge in formal languages, to capitalize on the complementary strengths of subsymbolic pattern recognition and symbolic compositionality [bhuyan2024neuro, nawaz2025review].
+Within knowledge-graph research, most neuro-symbolic work addresses link prediction or graph completion, where the hybrid models offer interpretable rules, guided learning, and more transparent decision pathways [delong2024neurosymbolic].
 Yet, like standard supervised techniques, many of these solutions still demand sizable labeled datasets.
@@
 Particle Filter and Statistical Reasoning.
-Particle filters–sequential Monte Carlo algorithms for recursive Bayesian estimation–propagate a set of weighted samples to approximate the posterior state distribution, outperforming classical Kalman filters in strongly non-linear or non-Gaussian settings (Künsch 2013).
-They are widely used in surveillance, robotics, and navigation, leveraging visual, geometric, and motion cues such as color, texture, and shape (Awal et al.
-2023).
+Particle filter is a sequential Monte Carlo algorithms for recursive Bayesian estimation [chen2003bayesian], it propagates a set of weighted samples to approximate the posterior state distribution, outperforming classical Kalman filters in strongly non-linear or non-Gaussian settings [kunsch2013particle, 5985520].
+They are widely used in surveillance, robotics, and navigation, leveraging visual, geometric, and motion cues such as color, texture, and shape [awal2023particle].
 In CADAR, we model graph nodes and edges as stochastic particles, enabling robust reasoning under noisy or inconsistent VLM outputs.
@@
 4), which reasons over the particle filter to identify cognitive attacks.
-Figure 2: CADAR system overview.
-Sequential video frames are converted into temporal perception graphs, which the attack-detection module then analyzes to identify, classify, and localize adversarial attacks.
+Figure 2: A system overview of CADAR.
+Sequential video frames are first transformed into spatial-temporal symbolic perception graphs.
+The attack-detection module then analyzes these graphs to identify, classify, and localize adversarial attacks, such as the visual modification and injection attacks shown in the examples.
 3.1 Cognitive Attacks
-We study four representative adversarial attack types that compromise an AR scene’s semantics (Fig. 1).
-Text-modification Attack.
-Altering on-scene text, e.g., overlaying “Free Parking” atop a “No Parking” sign, reverses instructions or injects false information, misleading both users and perception systems.
-Visual-modification Attack.
-Altering an object’s appearance or position (green light to red, stop sign moved to mid-junction) distorts recognition and user judgment.
-Removal Attack.
-Occluding or deleting safety-critical cues (e.g., an exit sign) erases vital information and breaks expected graph relations.
-Addition Attack.
-Inserting fictitious objects or labels (e.g., a fake hazard symbol) plants misleading cues that divert attention and corrupt downstream reasoning.
+We investigate four representative cognitive attack types that deliberately compromise the semantic integrity of AR scenes (Fig.1).
+The text-modification attack manipulates on-scene text—such as changing “No Parking” to “Free Parking”—to reverse intended meanings or inject false instructions, thereby deceiving both human users and perception systems.
+The visual-modification attack distorts object appearance or placement—for example, turning a green light to red or relocating a stop sign—leading to recognition errors and misinformed user decisions.
+The obstruction attack targets critical information by occluding or deleting essential cues like exit signs, disrupting safety awareness and expected graph relations.
+Finally, the injection attack introduces fictitious elements, such as fake hazard symbols or virtual labels, that embed misleading cues, divert user attention, and corrupt subsequent reasoning processes.
 3.2 Perception Graph Model
-Figure 3: Perception Graph generation: Given a video frame and its contextual description, generate the corresponding perception graph at time step t.
+Figure 3: Symbolic perception graph generation: given a video frame and its contextual description, VLMs generate the corresponding perception graph at time step t.
 The perception graph model is a symbolic representation of perception, simulating how humans perceive, interpret, and prioritize information.
@@
 Objects are represented as nodes, and object relationships are represented as edges.
-Given the context information and prompt (shown in the Figure  3), the VLM would extract the related objects with their relationships in the scene.
+Given the context information and prompt (shown in the Fig.3), the VLM would extract the related objects with their relationships in the scene.
 At timestep $t$, the Perception Graph $G_{t}$ is generated as follows.
-Graph Construction.
+3.2.1 Graph Construction.
 At discrete timestep $t\!\in\!\{1,\dots,T\}$ we ingest
@@
 $\hat{G}_{t}=(\mathcal{V}_{1:t},\,\mathcal{E}_{1:t}),\quad\mathcal{V}_{1:t}=\bigcup_{s=1}^{t}\mathcal{V}_{s},\quad\mathcal{E}_{1:t}=\bigcup_{s=1}^{t}\mathcal{E}_{s}.$ (4)
-Thus the GraphSet $\hat{G}_{t}$ is the union of all spatial and temporal connections that stitch object instances across time steps 1:t1{:}t.
+Thus, the GraphSet $\hat{G}_{t}$ is the union of all spatial and temporal connections that stitch object instances across time steps 1:t1{:}t.
 Each node still carries its original timestamp to preserve the temporal ordering.
@@
 Edge importance is derived from the contextual salience of the connected nodes.
-NodeSet and EdgeSet.
+3.2.2 NodeSet and EdgeSet.
 After defining the temporal graph structure $\hat{G}_{t}$, which links perception graphs across time, we introduce the NodeSet and EdgeSet to represent the temporally connected nodes and edges, respectively.
@@
 Thus the GraphSet $\hat{G}_{t}$ can also represented by NodeSet and EdgeSet:
-$\hat{G}_{t}=\bigl{(}\mathcal{V}_{1:t},\,\mathcal{E}_{1:t}\bigr{)},\quad\mathcal{V}_{1:t}=\!\!\bigcup_{i\in\mathcal{I}_{t}}\!\!\mathcal{N}_{i,t},\quad E_{1:t}=\!\!\bigcup_{i,j\in\mathcal{I}_{t}}\!\!\mathcal{E}_{ij,t}.$ (7)
+$\hat{G}_{t}=\bigl(\mathcal{V}_{1:t},\,\mathcal{E}_{1:t}\bigr),\hskip 2.84526pt\mathcal{V}_{1:t}=\!\!\bigcup_{i\in\mathcal{I}_{t}}\!\!\mathcal{N}_{i,t},\hskip 2.84526ptE_{1:t}=\!\!\bigcup_{i,j\in\mathcal{I}_{t}}\!\!\mathcal{E}_{ij,t}.$ (7)
 3.3 Particle Filter Model
 Our particle filter model tracks graph evolution and analyzes adversarial perturbations.
-It proceeds in three stages: prediction and matching, attack detection, and state estimation (cf.
-Fig.
+It proceeds in three stages: prediction and matching, attack detection, and state estimation (Fig.
 4).
@@
 The attack detection algorithm applies statistical tests to newly matched particles and discards those marked as attacked.
-Finally, weighting, resampling, and estimation produces an updated, temporally consistent graph state, preserving persistent NodeSet/EdgeSet identities and the relational structure over time.
+Finally, weighting, resampling, and estimation produce an updated, temporally consistent graph state, preserving persistent NodeSet/EdgeSet identities and the relational structure over time.
 Figure 4: The particle filter-based attack detection framework consists of three main modules: matching, attack detection, and state estimation.
-Prediction and Matching.
+3.3.1 Prediction and Matching.
 To maintain persistent object identities across frames, we equip the NodeSet with the Prediction-Matching algorithm and use it to decide whether a newly detected Node belongs to an existing object list or creates a new one.
@@
 To match the observed node to NodeSet, given a fresh node $v_{j}^{t}$ and a candidate NodeSet $i$, define
-$s_{j,i,m}\;=\;\frac{1}{|\mathcal{N}_{i,t-1}|}\sum_{v\in\mathcal{N}_{i,t-1}}\operatorname{sim}_{m}(v_{j}^{t},v),\qquad m\in\{n,d,f\}.$ (9)
+$s_{j,i,m}\;=\;\frac{1}{|\mathcal{N}_{i,t-1}|}\sum_{v\in\mathcal{N}_{i,t-1}}\operatorname{sim}_{m}(v_{j}^{t},v),\quad m\in\{n,d,f\}.$ (9)
 Candidate filter ($\leq 1$std in every modality):
 $\mathcal{C}_{j}\;=\;\Bigl\{i\in\mathcal{A}_{t-1}\;\Bigl|\;\bigl|s_{j,i,m}-\mu_{i,m}^{\,t-1}\bigr|\leq\sigma_{i,m}^{\,t-1},\;\forall m\in\{n,d,f\}\Bigr\}.$ (10)
-For each remaining candidate, measure the distance
-$D_{j,i}\;=\;\sum_{m\in\{n,d,f\}}\bigl{|}s_{j,i,m}-\mu_{i,m}^{\,t-1}\bigr{|}.$ (11)
-Finally, the assignment rule:
-$\mu_{t}(v_{j,t})=\begin{cases}\displaystyle\arg\min_{i\in\mathcal{C}_{j}}D_{j,i},&\mathcal{C}_{j}\neq\varnothing,\\[6.0pt] \text{new\_id},&\mathcal{C}_{j}=\varnothing.\end{cases}$ (12)
-If $\mu_{t}(v_{j}^{t})$ returns an existing ID, append $v_{j}^{t}$ to that NodeSet and update $A_{i}^{t}$; otherwise spawn a new NodeSet.
+Finally, the assignment rule based on measuring the remaining candidate distance:
+$\mu_{t}(v_{j,t})=\begin{cases}\displaystyle\arg\min_{i\in\mathcal{C}_{j}}\sum_{m\in\{n,d,f\}}\bigl|s_{j,i,m}-\mu_{i,m}^{\,t-1}\bigr|,&\mathcal{C}_{j}\neq\varnothing,\\[6.0pt] \text{new\_id},&\mathcal{C}_{j}=\varnothing.\end{cases}$ (11)
+If $\mu_{t}(v_{j}^{t})$ returns an existing ID, append $v_{j}^{t}$ to that NodeSet and update $A_{i}^{t}$; otherwise, spawn a new NodeSet.
 Confirmed edges $e_{ab}^{t}$ are then inserted into $E_{\mu_{t}(a)\,\mu_{t}(b),t}$ as before.
 After matching, the NodeSet goes into the Attack detection process.
-Attack Detection.
+3.3.2 Attack Detection.
 To detect text (Type 1) and visual (Type 2) modification attacks, those that manifest as anomalies in embedding space, we employ a dedicated statistical procedure.
@@
 $\mathbf{W}_{i,m}^{\,t-1}\;=\;\operatorname{PCA}\!\bigl(\{\mathbf{m}(v)\}_{v\in\tilde{\mathcal{N}}_{i,t-1}},\;k_{m}\bigr),$ (12)
-where $k_{m}$ is chosen such that $\geq\!99\%$ variance is retained. The projected embeddings are
-$\mathbf{z}_{m}(v)=\mathbf{W}_{i,m}^{\,t-1\,\top}\mathbf{m}(v)\;\in\mathbb{R}^{k_{m}}.$ (14)
-For each modality, fit the Gaussian reference model
-$\mathbf{z}_{m}(v)\;\sim\;\mathcal{N}\!\bigl{(}\mathbf{c}_{i,m}^{\,t-1},\,\boldsymbol{\Sigma}_{i,m}^{\,t-1}\bigr{)},\quad v\in\tilde{\mathcal{N}}_{i,t-1},$ (15)
+where $k_{m}$ is chosen such that $\geq\!99\%$ variance is retained. the resulting embeddings are modeled as Gaussian distributions:
+$\mathbf{z}_{m}(v)=\mathbf{W}_{i,m}^{\,t-1\,\top}\mathbf{m}(v)\;\sim\;\mathcal{N}\!\bigl(\mathbf{c}_{i,m}^{\,t-1},\,\boldsymbol{\Sigma}_{i,m}^{\,t-1}\bigr)$ (13)
 with the centroid and covariance of PCA-reduced embeddings
-$\mathbf{c}_{i,m}^{\,t-1}=\frac{1}{|\tilde{\mathcal{N}}_{i,t-1}|}\sum_{v\in\tilde{\mathcal{N}}_{i,t-1}}\mathbf{z}_{m}(v),$ (16)
-$\boldsymbol{\Sigma}_{i,m}^{\,t-1}=\mathrm{Cov}\!\bigl{\{}\mathbf{z}_{m}(v)\bigr{\}}.$ (17)
+$\mathbf{c}_{i,m}^{\,t-1}=\frac{1}{|\tilde{\mathcal{N}}_{i,t-1}|}\sum_{v\in\tilde{\mathcal{N}}_{i,t-1}}\mathbf{z}_{m}(v),\hskip 2.84526pt\boldsymbol{\Sigma}_{i,m}^{\,t-1}=\mathrm{Cov}\!\bigl\{\mathbf{z}_{m}(v)\bigr\}.$ (14)
 Use the Mahalanobis distance for measuring Text and Visual Modification attack. For a new node $v_{i}^{t}$ compute
@@
 $v_{i}^{t}\;\text{as}\;\begin{cases}\text{Type 1 Attack(text)}&\text{if }\mathcal{M}_{i,t}^{(d)}>\chi^{2}_{k_{d},\alpha},\\[2.0pt] \text{Type 2 Attack(image)}&\text{if }\mathcal{M}_{i,t}^{(f)}>\chi^{2}_{k_{f},\alpha}.\end{cases}$ (16)
-For the structural attacks, specifically object removal (Type 3) and object addition (Type 4), the detector relies on lightweight temporal heuristics rather than statistical distance.
-After each matching step the algorithm keeps a two-frame sliding window of NodeSet activity.
-If an object that previously carried a high task-importance score ($\pi\geq\pi_{\text{high}}$) fails to obtain a match in both the current and the immediately preceding frame, the system labels the disappearance a removal attack.
-Conversely, any NodeSet that first appears within the last two frames and whose reasonability score falls below a predefined threshold ($\rho\leq\rho_{\text{low}}$) is treated as an addition attack.
+For the structural attacks, specifically object obstruction (Type 3) and object injection (Type 4), the detector relies on lightweight temporal heuristics rather than statistical distance.
+After each matching step, the algorithm keeps a two-frame sliding window of NodeSet activity.
+If an object that previously carried a high task-importance score ($\pi\geq\pi_{\text{high}}$) fails to obtain a match in both the current and the immediately preceding frame, the system labels the disappearance an obstruction attack.
+Conversely, any NodeSet that first appears within the last two frames and whose reasonability score falls below a predefined threshold ($\rho\leq\rho_{\text{low}}$) is treated as an injection attack.
 In the attack detection final step, if $v_{i}^{t}$ is flagged as attack, it is appended to the $\mathcal{N}_{i,t}$ but excluded from future Gaussian updates; otherwise going to the Node Estimation process, including weight and resample, then stored in a clean reference $\tilde{\mathcal{N}}_{i,t}$.
-Estimation.
+3.3.3 Estimation.
 After a new observation $v_{i,t}$ has been matched to object $i$ and detected as a non-attack node, the particle filter estimates the new node by refining that observation by resampling historical detections that resemble it.
@@
 So $w_{j}$ is the normalized weight of the particle, based on its distance to the new particle.
-Then use the weight to resample $M$ indices $s_{1},\dots,s_{M}\sim\operatorname{Multinomial}(w_{1:M})$ yields
-$\hat{a}\;=\;\frac{1}{M}\sum_{k=1}^{M}a^{(s_{k})},$ (22)
+Similarly, for embedding attributes, modality $m\in\{n,d,f\}$ (name, description, image) let ${m}^{\star}={m}(v_{i}^{t})$ and ${m}^{(j)}={m}(v^{(j)})$.
+Weight is defined as
+$sim_{j}=\frac{{m}^{(j)}\!\cdot\!{m}^{\star}}{\|{m}^{(j)}\|\,\|{m}^{\star}\|},\qquad w_{j}=\frac{\exp(\lambda_{m}sim_{j})}{\sum_{k=1}^{M}\exp(\lambda_{m}sim_{k})},$ (19)
+with a scale factor $\lambda_{m}>0$.
+Then use the weight to resample $M$ indices $s_{1},\dots,s_{M}\sim\operatorname{Multinomial}(w_{1:M})$ yields estimated scalar particle and embedding particle:
+$\hat{a}\;=\;\frac{1}{M}\sum_{k=1}^{M}a^{(s_{k})},\quad\hat{{m}}\;=\;\frac{\sum_{k=1}^{M}{m}^{(s_{k})}}{\Bigl\|\sum_{k=1}^{M}{m}^{(s_{k})}\Bigr\|},$ (20)
 where $\hat{a}$ is the estimated new scalar particle.
-For embedding attributes, modality $m\in\{n,d,f\}$ (name, description, image) let $\mathbf{m}^{\star}=\mathbf{m}(v_{i}^{t})$ and $\mathbf{m}^{(j)}=\mathbf{m}(v^{(j)})$. Cosine similarity gives
-$sim_{j}=\frac{\mathbf{m}^{(j)}\!\cdot\!\mathbf{m}^{\star}}{\|\mathbf{m}^{(j)}\|\,\|\mathbf{m}^{\star}\|},\qquad w_{j}=\frac{\exp(\lambda_{m}sim_{j})}{\sum_{k=1}^{M}\exp(\lambda_{m}sim_{k})},$ (23)
-with a scale factor $\lambda_{m}>0$. Resampling as above yields indices $s_{1},\dots,s_{M}\sim\operatorname{Multinomial}(w_{1:M})$ and the unit-normalised mean
-$\hat{\mathbf{m}}\;=\;\frac{\sum_{k=1}^{M}\mathbf{m}^{(s_{k})}}{\Bigl{\|}\sum_{k=1}^{M}\mathbf{m}^{(s_{k})}\Bigr{\|}},$ (24)
-would be the resampled embedding particle.
 Finally, replacing each raw attribute in $v_{i}^{t}$ by its resampled estimate gives the estimated node
-$\hat{v}_{i,t}=\bigl{(}\hat{p},\;\hat{\pi},\;\hat{\rho},\;\hat{\mathbf{n}},\;\hat{\mathbf{d}},\;\hat{\mathbf{f}}\bigr{)}$, (25)
-which becomes the particle filter’s estimation state for object $i$ at time $t$. If $M<5$ (insufficient history) we skip resampling and simply set $\hat{v}_{i,t}=v_{i,t}$.
+[…]
+which becomes the particle filter’s estimation state for object $i$ at time $t$. If $M<5$ (insufficient history), we skip resampling and simply set $\hat{v}_{i,t}=v_{i,t}$.
 4 Experiment
-Table 1: Results of the state-of-the-art methods and CADAR on Extended AR-VIM Dataset
-Method Attack Detection: Accuracy/F1-score (%) Localization Acc. of attack(%)
-Text Visual Removal Addition Text+Visual Add.+Rem. Overall
-CADAR (Ours) 80.2/ 74.9 82.3/ 76.1 83.0/ 86.3 82.1/ 83.6 67.0/ 69.7 86.1/ 75.2 80.1/77.6 68.0
-GPT-o3 62.5/56.6 23.8/29.4 76.2/66.7 61.9/55.3 16.7/25.0 14.3/25.0 39.2/39.8 -
-Gemini 2.5 66.7/71.8 66.7/77.8 81.0/69.4 90.5/88.4 14.3/18.1 71.4/71.4 69.4/71.1 -
-3D ResNet 12.5/22.2 57.1/72.7 57.1/34.7 42.8/31.6 - - 42.4/40.3 -
-ViViT 50.0/44.4 28.6/33.3 71.4/47.6 14.3/20.0 - - 41.1/36.3 -
-In this section, we provide a detailed description of the CADAR implementation and the experimental settings used for our evaluation.
-We then present and analyze the numerical results of our experiments, demonstrating that CADAR significantly outperforms several baselines, achieving an overall accuracy of 80.1% compared to 69.4% for the next-best model.
+Table 1: Attack-wise Accuracy and F1-score (%) across scenes for frame-level and video-level baselines.
+Each attack type forms a block, with sub-rows indicating methods.
+Columns represent scene categories.
+CADAR outperforms VLM-based (GPT-5-mini, Gemini-2.5-flash) and supervised video models (3D-ResNet, ViViT) in 23 out of 25 attack–scene blocks.
+Class-specific Accuracy / F1-score (%)
+Attack Detection Method School&Office Traffic&Mobility Industrial&Infra Medical&Lab Public&Home Avg Acc/F1
+[…]
+GPT-5-mini 50.6 / 67.2 59.6 / 74.7 52.1 / 68.5 52.5 / 68.9 61.1 / 75.9 55.2 / 71.0
+Gemini-2.5-flash 73.0 / 84.4 58.4 / 73.7 77.4 / 87.2 41.6 / 58.8 51.4 / 67.9 60.4 / 74.4
+Video 3D ResNet 12.5 / 22.2 10.9 / 19.6 12.8 / 22.6 15.6 / 27.0 40.0 / 57.1 18.4 / 29.7
+ViViT 26.3 / 41.7 70.6 / 82.8 25.0 / 42.1 33.3 / 50.0 43.5 / 60.6 39.7 / 55.4
+[…]
+GPT-5-mini 51.0 / 67.6 61.5 / 76.2 50.6 / 67.2 45.5 / 62.5 59.8 / 74.9 53.7 / 69.7
+Gemini-2.5-flash 49.5 / 66.3 76.7 / 86.8 43.9 / 61.0 58.7 / 74.0 40.8 / 58.0 53.9 / 69.2
+Video 3D ResNet 16.5 / 28.3 7.2 / 13.3 8.3 / 15.3 12.7 / 22.5 8.3 / 15.4 10.6 / 19.0
+ViViT 22.2 / 36.4 44.4 / 61.5 20.0 / 33.3 28.6 / 44.4 25.0 / 40.0 28.0 / 43.1
+[…]
+GPT-5-mini 75.4 / 85.9 46.0 / 63.0 61.7 / 76.3 46.3 / 63.3 63.2 / 77.4 58.5 / 73.2
+Gemini-2.5-flash 64.7 / 78.6 51.6 / 68.1 55.0 / 71.0 36.9 / 53.9 43.6 / 60.7 50.4 / 66.5
+Video 3D ResNet 8.1 / 15.0 3.0 / 5.9 5.0 / 9.5 9.4 / 17.2 4.5 / 8.7 6.0 / 11.3
+ViViT 50.0 / 66.7 54.5 / 70.6 46.2 / 63.2 71.4 / 83.3 41.7 / 42.3 52.8 / 65.2
+[…]
+GPT-5-mini 69.1 / 81.7 60.7 / 75.6 56.3 / 72.0 55.2 / 71.2 63.3 / 77.5 60.9 / 75.6
+Gemini-2.5-flash 70.1 / 82.4 54.1 / 70.2 62.8 / 77.2 61.5 / 76.1 53.1 / 69.4 60.3 / 75.0
+Video 3D ResNet 6.6 / 12.3 12.8 / 22.7 9.8 / 17.8 13.3 / 23.4 15.7 / 27.2 11.6 / 20.7
+ViViT 30.0 / 46.2 45.5 / 62.5 20.0 / 33.3 25.0 / 40.0 55.6 / 71.4 35.2 / 50.7
+[…]
+GPT-5-mini 71.8 / 83.6 52.1 / 68.5 62.8 / 77.1 70.5 / 82.7 68.1 / 81.0 65.1 / 77.4
+Gemini-2.5-flash 54.2 / 70.3 38.2 / 55.3 44.0 / 61.1 41.1 / 58.3 44.1 / 61.2 44.3 / 61.2
+Video 3D ResNet 8.5 / 15.7 11.7 / 20.9 8.6 / 15.9 15.9 / 27.5 29.4 / 45.5 14.8 / 25.1
+ViViT 40.0 / 57.1 15.4 / 26.7 7.7 / 14.3 40.0 / 57.1 30.8 / 47.1 26.8 / 40.5
+Overall Mean (per method): CADAR: 74.6 / 84.2   GPT-5-mini: 58.7 / 73.4   Gemini-2.5-flash: 53.9 / 69.3   3D ResNet: 12.3 / 21.1   ViViT: 36.5 / 51.0
+Table 2: Recall comparison for hybrid-attack scenarios(%), CADAR achieves the highest recall across both hybrid combinations.
+Hybrid Attack Type Recall of Method (%)
+GPT-5-mini Grok-4-fast CADAR Gemini-2.5-flash
+Text+Visual 35.0 41.2 76.4 29.9
+Obstruction+Injection 32.7 26.1 80.4 45.8
+Table 3: Recall (%) of attack detection under different numbers of simultaneous attacks.
+CADAR maintains strong performance as the number of attacks increases, whereas baseline models degrade substantially.
+Method Number of Attacks Avg
+1 Attack 2 Attacks 3 Attacks 4 Attacks
+GPT5-mini 66.4 40.2 31.5 29.6 41.9
+Grok-4-fast 69.7 55.0 51.3 36.5 53.1
+Gemini-2.5-flash 61.7 43.6 33.0 31.5 42.5
+[…]
+In this section, we provide a detailed description of the CADAR experimental settings used for performance evaluation.
+We then present and analyze the numerical results of our experiments, demonstrating that CADAR significantly outperforms several baselines.
 An ablation study is also conducted to analyze the contributions of our key components.
-4.1 Implementation of CADAR
-CADAR is implemented using a modular architecture.
-The perception graph is constructed and managed through several key components.
-The Node.py and Edge.py modules generate multimodal object representations by combining textual and visual features: GPT-4.1 is used for generating descriptive text, google/owlv2-base-patch16-ensemble (Minderer, Gritsenko, and Houlsby 2024) enables zero-shot object detection, and SAM2 (Ravi et al.
-2024) provides object segmentation.
-Embeddings are extracted using all-MiniLM-L6-v2 for object names, all-mpnet-base-v2 for descriptions, and ResNet-34 for image features.
-The Graph.py module manages the perception graph structure, maintaining node–edge relationships and temporal information.
-Temporal reasoning is handled by NodeSet.py and EdgeSet.py, which update the graph across frames, detect potential attacks, and evolve the graph state.
-Attack detection is based on statistical anomaly detection using PCA to preserve 99% of variance, with thresholds computed from a 99% confidence level of the chi-squared distribution.
-Finally, Utils.py provides supporting tools for video processing and mathematical computations.
 4.1 Experimental Settings
-Extended AR-VIM Dataset.
-We base our experiments on an extended version of the AR-VIM dataset, which is specifically designed to evaluate visual information manipulation in AR environments (Xiu and Gorlatova 2025).
-The Extended AR-VIM is the first publicly available AR cognitive attack dataset (details in Appendix).
-Each sample in the dataset consists of a short video collected by an AR device (Meta Quest 3 and Samsung Galaxy S24), approximately 5-10 seconds in length, where virtual content appears at a specific frame to trigger an attack.
-The dataset has 160 videos in different scenes, spanning all four cognitive-attack classes.
-The scene covers transportation (e.g., highway, train, aircraft), experimental settings (e.g., lab equipment), industrial environments (e.g., construction, agriculture), workspaces (e.g., warehouse, office), as well as medical, daily-life, and public-safety domains.
-These range from altering freeway exit-ramp signage to suppressing voltage warnings on live equipment.
-Together, these videos provide a diverse, realistic, and tightly controlled test bed for evaluating the perception-level of cognitive attacks in AR environments.
-The extended AR-VIM dataset consists of 41 removal attack samples, 35 addition attack samples, 40 text manipulation attack samples, 30 visual manipulation attack samples, and 14 mix-class attack (text + visual manipulation attacks and addition + removal attacks) samples.
-For baseline training models, it also contains 40 non-attacked videos.
-Considering the scene availability, the dataset was collected using two complementary methods: (1) The monitor-based method places AR content on a monitor screen with background images and captures it via video recording, allowing rapid and controlled generation of attack instances; (2) The real-world-based method, in contrast, uses AR headsets to capture first-person views in physical environments, introducing natural visual noise and depth information.
-Employing both methods is critical to capturing the trade-off between efficient data generation and real-world fidelity.
-Task and Metrics.
-We evaluate our model on the task of attack classification in videos.
-CADAR supports frame-wise detection, variable-length multi-label outputs, and localization of multiple simultaneous attacks, which are capabilities that few existing models possess.
-As most baselines can only classify attacks at the video level and lack the capability for fine-grained, frame-level detection or localization, we compare CADAR against both advanced reasoning VLMs (GPT-o3 and Google Gemini 2.5 (Comanici et al.
-2025)) and trained video classification models, including 3D-ResNet (Tran et al.
-2018) and ViViT (Arnab et al.
-2021).
-We use accuracy and F1-score to evaluate attack detection performance: frame-wise for CADAR, and video-level for the baselines.
-Additionally, we report attack localization accuracy, which measures how precisely the model locates the attacks in the frame.
-Results are summarized in Table 1.
-For mixed-attack instances, a detection is considered correct only if both attack types are correctly identified.
+CADAR-50K Dataset.
+The CADAR-50K dataset extends the AR-VIM dataset [11202433], which was designed to evaluate visual information manipulation in AR environments.
+CADAR-50K is the first publicly available dataset for AR cognitive-attack analysis, containing 375 videos across diverse scenes and more than 50K frames, covering all seven categories (Fig.
+5 and Fig.
+6).
+Each sample is an 8–12 second AR-captured video (recorded using a Meta Quest 3 or Samsung Galaxy S24), where virtual content is introduced at a specific frame to induce an attack.
+For the non-attack elements, we use irrelevant background objects, benign AR elements, and semantically benign prompts that can interfere with VLM outputs but do not alter the human-perceived task semantics.
+Considering scene availability, the dataset was collected using two complementary methods: (1) a virtual-scene method, where fully synthetic environments are first generated and displayed on a monitor or VR headset, and AR content is then overlaid onto these virtual scenes before being captured through video recording.
+This enables rapid, controllable creation of diverse attack scenarios; and(2) a real-world method, where AR headsets record first-person views in physical environments, introducing natural noise, lighting variation, and depth cues.
+Task, Baselines and Metrics.
+Figure 5: CADAR-50K dataset overview.
+We evaluate our model on the task of video-based cognitive attack classification.
+The CADAR dataset supports frame-wise detection, variable-length multi-label outputs, and localization of simultaneous attacks.
+We compare CADAR against two categories of baselines that similarly accept multiple reference frames as input and output attack classes: (1) multimodal VLMs, including state-of-the-art OpenAI GPT-5, Google Gemini-2.5 [comanici2025gemini] and xAI Grok-4, which perform fine-grained frame-level detection; and (2) supervised video models, the CNN-based 3D-ResNet [tran2018closer] and transformer-based ViViT [arnab2021vivit], which output a single prediction per video.
+In our first experiment(Table 1), we use class-specific accuracy and F1-score to assess overall detection performance.
+The class-specific accuracy is defined as $\mathrm{TP}/(\mathrm{TP}+\mathrm{FN}+\mathrm{FP})$, which penalizes both missed detections and false alarms for that class while ignoring true negatives that dominate under severe class imbalance.
+For hybrid attack or multiple attack videos(Table 2 and 3), a prediction is considered correct only if all attack types are simultaneously identified.
+We also report per-class recall , $\mathrm{TP}/(\mathrm{TP}+\mathrm{FN})$, to measure correctness conditioned on the presence of the class.
 4.2 Numerical Results and discussion
-Table 1 presents the evaluation results for attack classification on the extended AR-VIM dataset.
-CADAR outperforms both the supervised learning baselines and the pre-trained VLMs.
-Among the different attack types, removal and addition attacks are generally easier to detect, leading to higher accuracy compared to text and visual modifications.
-Specifically, CADAR achieves a detection accuracy of 83.0% for removal attacks and 82.1% for addition attacks, compared to 80.2% for text and 82.3% for visual attacks.
-Notably, CADAR demonstrates balanced and robust performance across all attack categories, achieving superior accuracy and F1 scores.
-For instance, its overall accuracy/F1 score is 80.1%/77.6%, significantly higher than the next best model, Gemini 2.5, which scores 69.4%/71.1%.
-Supervised models such as 3D-ResNet and ViViT require large-scale training data and perform poorly in zero- or few-shot scenarios.
-For example, 3D-ResNet achieves an overall accuracy of 42.4%, and ViViT achieves 41.1%.
-In contrast, pre-trained VLMs leverage prior knowledge and exhibit strong reasoning abilities, but their outputs are often unstable and prone to hallucination.
-For example, GPT-o3 and Gemini 2.5 achieve overall accuracies of 39.2% and 69.4%, respectively.
-CADAR offers a novel approach that addresses these limitations and achieves the best overall performance.
-However, it is not without shortcomings.
-The localization accuracy is constrained by the capabilities of the zero-shot object detection and segmentation models, which struggle in complex scenes or with unfamiliar objects.
-CADAR achieves a localization accuracy of 68.0%.
-Moreover, reduced performance on text attacks suggests that highly subtle edits, e.g., a single-character change that scarcely alters semantics or appearance, remain difficult for the detector.
+Cross-Scene Cognitive Attack Classification.
+The results in Table 1 demonstrate that CADAR achieves the highest overall accuracy(74.6%) and F1-score(84.2%) and its average accuracy exceeding the next best baseline models (GPT-5-mini) by 16%.
+CADAR attains the highest class-specific accuracy and F1 in 23 out of 25 attack–scene blocks.
+The margin is especially large for Visual and Obstruction attacks, where CADAR delivers strong gains, underscoring its robustness to fine-grained semantic distortions.
+We also observe that VLM baselines exhibit substantial accuracy fluctuations across different scenes, indicating that their reasoning processes are unstable and prone to hallucination.
+Because they rely on jointly processing all textual and visual inputs, they become increasingly sensitive to misleading cues, malicious prompts, and complex visual environments.
+This vulnerability highlights the advantage of our neuro-symbolic approach, which offers more controlled and explainable reasoning under diverse AR scenarios.
+On the other hand, traditional video-classification models (3D-ResNet and ViViT) perform poorly across all attack types, suggesting that the models are inadequate for capturing the fine-grained semantic inconsistencies central to cognitive attack detection.
+Their limited performance is further exacerbated by the lack of large-scale, AR-specific training data, making these models particularly weak in zero- or few-shot scenarios.
+Overall, the results indicate that CADAR’s neurosymbolic design and temporal reasoning components provide strong advantages for detecting subtle, localized, and multi-instance AR cognitive attacks, achieving robust performance across diverse environments and attack conditions.
+Figure 6: Visualization frame examples of CADAR outputs across five scene categories in CADAR-50K 375 videos.
+Rows show the Reference Scene, Detection Scene, and CADAR Output.
+Output attack types: I = Injection, V = Visual, T = Text, O = Obstruction, N = Non-attack.
+Some examples contain hybrid or multiple attacks.
+Hybrid and Multiple Attacks
+Table 2 and 3 evaluate model robustness in hybrid and multi-attack settings.
+In hybrid-attack scenarios, CADAR attains the highest recall for both Text+Visual(76.4%) and Obstruction+Injection(80.4%) cases and surpasses all VLM baselines by a substantial margin.
+This demonstrates that the proposed neurosymbolic graph reasoning framework can effectively detect the attacks separately.
+In multiple attack scenarios, the performance of VLM baselines declines sharply as attack complexity grows, whereas CADAR retains strong recall, from 81.0% for a single attack to 67.2% for four simultaneous attacks, showing the smallest performance drop among all compared methods.
+These results highlight CADAR’s resilience in multi-attack scenarios and its clear advantage in detecting objects in parallel.
+Effect of Scene Complexity.
+The recall trends in Fig.
+7(a) reveal a clear divergence between VLM-based baselines and CADAR as scene complexity increases.
+For GPT-5-mini and Gemini-2.5-flash, recall drops sharply when the number of key objects and frames increases, suggesting that conventional VLMs struggle with reasoning over long visual and textual context and are easily affected by irrelevant or distracting content.
+In contrast, CADAR performance remains stable as reference frames increase, since it extracts complex scene information into a structural spatial-temporal perception graph.
+CADAR improves its decision accuracy by accumulating prior evidence through additional reference frames, and evaluates objects and attacks separately rather than relying on a single holistic reasoning step.
+The CADAR design results in substantially smaller performance degradation compared with standard VLMs.
 4.3 Ablation Study
 Estimation Module.
-Figure 5: Ablation study on estimation module and reference frames.
+Figure 7: Recall under increasing reference frames and object quantities (a), and ablation studies on the estimation module (b) and the number of reference frames (c).
 Fig.
-5(a) shows our ablation study examining the impact of the estimation module in particle filtering.
+7(b) shows our ablation study examining the impact of the estimation module in particle filtering.
 When a polluted node that might have bad data due to VLM unstable output or due to an undetected attack, the estimation module’s weighting and resampling can effectively prevent it from corrupting the ground truth knowledge.
-In the ablation study, we choose 12 videos for all attack types.
-As shown in the left plot of the figure, for text and visual modification attacks, the detection accuracy drops significantly as the number of frames without the estimation module increases.
-Specifically, the accuracy for text attacks decreases from approximately $83\%$ to $69\%$ and from approximately $81\%$ to $55\%$ for visual attacks.
+We randomly sample 6 videos per base attack type (24 in total) from the test split for the ablation study.
+As shown in the left plot of the figure, for text and visual modification attacks, the recall drops significantly as the number of frames without the estimation module increases.
+Specifically, the recall for text attacks decreases from approximately $81\%$ to $63\%$ and from approximately $79\%$ to $55\%$ for visual attacks.
 This indicates that the estimation module is crucial for correctly identifying and mitigating errors in more subtle attack types.
-In contrast, for removal and addition attacks, the accuracy remains relatively stable, suggesting these attacks are more straightforward to detect and less dependent on the estimation module’s corrective capabilities.
+In contrast, for obstruction and injection attacks, the recall remains relatively stable, suggesting these attacks are more straightforward to detect and less dependent on the estimation module’s corrective capabilities.
 Reference Frames.
 Fig.
-5(b) shows the impact of the number of reference frames on detection accuracy.
-A larger number of reference frames, or more prior knowledge of the scene, will result in higher accuracy.
+7(c) shows the impact of the number of reference frames on recall.
+A larger number of reference frames, or more prior knowledge of the scene, will result in higher recall.
 For text and visual modification attacks, increasing the number of reference frames from 3 to 11 significantly improves performance.
-The accuracy for text attacks increases from around $55\%$ to $80\%$ and from around $50\%$ to $78\%$ for visual attacks.
+The recall for text attacks increases from around $52\%$ to $83\%$ and from around $48\%$ to $79\%$ for visual attacks.
 This demonstrates that a larger set of reference frames provides more prior knowledge for more accurate statistical reasoning, which is essential for detecting subtle text and visual changes.
-Similar to the estimation module study, the detection accuracy for removal and addition attacks remains largely unaffected by the number of reference frames, consistently staying above $80\%$.
+Similar to the estimation module study, the recall for obstruction and injection attacks remains largely unaffected by the number of reference frames, consistently staying around $80\%$.
 This reinforces the finding that these attacks are less sensitive to the amount of contextual information and prior knowledge.
@@
 We presented CADAR, a neuro-symbolic framework that couples a perception graph with particle-filter reasoning to yield interpretable attack detection in AR/MR.
-On the Extended AR-VIM dataset, CADAR performs strongly across all attack types, and ablations highlight the value of our estimation module and richer reference histories.
-Limitations stem from the underlying open-vocabulary detection/segmentation, and miss very fine-grained augmentations.
-Future work will pursue task-specific fine-tuning of perception models, and larger-scale real-world evaluations and user studies.
+On the CADAR-50K dataset, CADAR performs strongly across all attack types, and ablations highlight the value of our estimation module and richer reference histories.
 Overall, CADAR highlights the value of neuro-symbolic design by combining the adaptability of foundation models with symbolic reasoning for a robust and interpretable AR safety mechanism.
-Appendix A Extended AR-VIM Dataset Samples
-Representative sample videos for each attack class (excluding the No Attack category) is shown below. The full dataset consists of 160 videos across diverse scenes.
-Figure 6: Sample dataset content for each AR Attack Class.
-Appendix B Dataset Construction
-For the real-world scenes in the Extended AR-VIM dataset, we collected additional samples using a Meta Quest 3 headset. The data collection setup is shown below
+Appendix
+Appendix A CADAR Implementation Details
+This section provides the full implementation details of CADAR, covering perception-graph construction, multimodal embedding generation, temporal reasoning, and statistical attack detection.
+The system follows a modular architecture implemented across the Node.py, Edge.py, Graph.py, NodeSet.py, EdgeSet.py, and Utils.py modules in our codebase.
+A.1 Perception Graph Construction
+Object Detection and Segmentation.
+For each frame, CADAR extracts object proposals using two vision modules:
+  • Zero-shot object detection:
+    google/owlv2-base-patch16-ensemble [minderer2024scalingopenvocabularyobjectdetection], which produces category-agnostic bounding boxes and confidence scores.
+  • Segmentation masks: SAM2 [ravi2024sam2segmentimages], used to obtain object-level silhouettes for robust visual embedding extraction.
+Textual Attribute Generation.
+To obtain semantic descriptions associated with the detected region, we use a lightweight VLM (GPT-4.1) generates short, structured textual descriptions.
+These descriptions include attributes such as object type, color, role, contextual relevance, and scene relationships.
+The attributes also include numerical importance, reasonability and probability.
+Multimodal Embedding Extraction.
+Each detected object is represented by three embedding modalities, each extracted using a specialized encoder:
+  • Name embedding: paraphrase-MiniLM-L6-v2 (384-d).
+  • Description embedding: all-mpnet-base-v2 (768-d).
+  • Visual embedding: a ResNet-34 feature encoder applied to SAM2-extracted object crops (512-d).
+These multimodal perceptual vectors are then projected using PCA for statistical reasoning (described below).
+Node and Edge Formation.
+Implementation Details.
+The software architecture is structured around two primary classes, Node.py and Edge.py, which manage the lifecycle and serialization of the perception graph components.
+The Node class orchestrates the bottom-up extraction pipeline, initializing the vision models to generate object masks and invoking the VLM to populate attribute fields (e.g., reasonability, importance) while computing and storing dense multimodal embeddings.
+Complementing this, the Edge class executes the relational analysis by ingesting the instantiated node list and source image; it prompts the LLM to infer pairwise dependencies (such as spatial or functional links) and encodes these descriptions into vector format, thereby establishing the graph’s connectivity before serializing the complete state to compressed binary files.
+A.2 Graph Backbone and Temporal Tracking
+Graph Management.
+The Graph.py module serves as the fundamental container for encapsulating the state of the perception system at discrete timestamps.
+It aggregates detected entities and their pairwise relationships into a coherent network structure, storing lists of Node and Edge instances alongside temporal metadata and references to the source imagery.
+Beyond acting as a static data structure, this module facilitates the serialization of the complete graph state into compressed binary formats to ensure efficient persistence and provides visualization utilities that render the probabilistic attributes and topological connectivity of the scene using network layout algorithms.
+NodeSet and EdgeSet Evolution.
+To transform isolated frame-level detections into continuous historical tracks, the system relies on the NodeSet.py and EdgeSet.py modules.
+The NodeSet class maintains the identity of objects over time by associating new detections with existing tracks via multi-modal embedding similarity (visual, semantic, and nominal) rather than simple spatial overlap.
+It implements a particle filter-based mechanism to estimate stable node attributes from noisy observations and employs statistical methods—such as PCA and Mahalanobis distance—to detect anomalies or adversarial attacks.
+Complementing this, the EdgeSet tracks the evolution of semantic and spatial relationships between these persistent entities, preserving the history of interactions to support high-level reasoning as the scene evolves.
+A.3 Statistical Anomaly Detection
+Dimensionality Reduction and Feature Whitening.
+To ensure robust distance metrics in high-dimensional spaces, we apply Principal Component Analysis (PCA) to the embedding history of each track.
+For a set of historical embedding vectors $\mathbf{X}\in\mathbb{R}^{N\times D}$, we compute a projection that preserves a specific variance ratio.
+Specifically, for node textual and visual embeddings, we retain 99% of the variance ($v=0.99$), while for edge description embeddings, we retain 90% ($v=0.90$).
+For visual embeddings, we additionally apply whitening to normalize the principal components to unit variance.
+This projection results in a dynamic latent dimension $k$ (typically $16\leq k\leq 32$), which defines the degrees of freedom for subsequent statistical tests.
+Regularized Mahalanobis Distance.
+We model the distribution of ”clean” reference frames for any given object or relationship as a multivariate Gaussian.
+To assess whether a new observation $x_{t}$ belongs to this distribution, we calculate the squared Mahalanobis distance.
+To prevent numerical instability when the covariance matrix is singular (common in short history windows), we introduce a regularization term $\epsilon\mathbf{I}$:
+$d^{2}(x_{t})=(x_{t}-\mu)^{\top}(\Sigma+\epsilon\mathbf{I})^{-1}(x_{t}-\mu)$ (22)
+where $\mu$ is the mean of the historical embeddings, $\Sigma$ is the covariance matrix, and $\epsilon=10^{-6}$ is a smoothing factor.
+Chi-Square Hypothesis Testing.
+Anomalies are identified by comparing the squared distance against a critical value derived from the Chi-square distribution with $k$ degrees of freedom.
+An observation is flagged as an attack (e.g., semantic modification or visual spoofing) if:
+$d^{2}(x_{t})>\chi^{2}_{1-\alpha,\,k}$ (23)
+where $\alpha$ represents the significance level.
+We employ adaptive sensitivity thresholds: $\alpha=0.10$ for textual deviations, $\alpha=0.05$ for visual deviations, and $\alpha=0.01$ for edge relationship deviations.
+This statistical framework allows CADAR to detect subtle adversarial manipulations that deviate from the established semantic baseline.
+A.4 Particle-Filter-Based Temporal Reasoning
+To mitigate the stochasticity of VLM outputs and detection noise, we implement a particle filter that maintains a belief state over node attributes.
+Unlike standard tracking which estimates position, our filter estimates semantic consistency.
+Weight Update.
+For a node $n$ with history $H$, we define the weight $w_{i}$ of the $i$-th particle (attribute hypothesis) based on its similarity to the incoming observation $z_{t}$.
+For scalar attributes (probability, importance, reasonability), weights follow a Gaussian kernel:
+$w_{i}\propto\exp\left(-\frac{(v_{i}-z_{t})^{2}}{2\sigma^{2}}\right)$ (24)
+For high-dimensional embedding attributes, weights are derived from cosine similarity, scaled by a confidence factor $\lambda$:
+$w_{i}\propto\exp\left(\lambda\cdot\frac{\mathbf{e}_{i}\cdot\mathbf{z}_{t}}{\|\mathbf{e}_{i}\|\|\mathbf{z}_{t}\|}\right)$ (25)
+where $\lambda=1$ is the scaling factor used to tune the confidence in the top matches.
+Resampling and Estimation.
+We perform importance resampling to filter outlier descriptions generated by the VLM.
+The final estimated state $\hat{x}_{t}$ is computed as the weighted centroid of the resampled population, effectively smoothing the semantic trajectory of the object and rejecting transient classification errors.
+Table 4: Hyper-parameters for Statistical Detection and Tracking.
+Parameter Value Description
+Matching Threshold 0.2 Similarity floor for node association
+Min. History Window 5 frames Minimum samples required for PCA/Mahalanobis
+PCA Variance (Nodes) 0.99 Variance retained for text/image embeddings
+PCA Variance (Edges) 0.90 Variance retained for edge descriptions
+Regularization ($\epsilon$) $10^{-6}$ Covariance matrix stabilization term
+$\alpha$ (Text) 0.10 Significance level for node text anomalies
+$\alpha$ (Image) 0.05 Significance level for visual appearance anomalies
+$\alpha$ (Edge) 0.01 Significance level for relationship anomalies
+Particle Scale ($\lambda$) 1.0 Exponential scaling for embedding similarity weights
+A.5 Supporting Utilities
+Utility and Pre-processing.
+The utils.py module serves as a comprehensive toolkit for low-level visual processing and mathematical operations underpinning the higher-level graph logic.
+A.6 Summary
+Overall, CADAR integrates zero-shot perception, multimodal symbolic representation, PCA-based statistical modeling, and particle-filter temporal reasoning into a unified framework for robust AR cognitive attack detection.
+The modular architecture allows flexible substitution of detectors, encoders, and reasoning mechanisms, while ensuring reproducibility across all experimental settings.
+Appendix B CADAR-50K Dataset Samples
+Representative sample videos for each class (excluding the No Attack category) is shown below. The full dataset consists of 375 videos across diverse scenes.
+Figure 8: Sample dataset content for each AR Attack Class. For video examples, please refer to CADAR_demo.mp4 in the supplementary materials.
+B.1 Dataset Construction
+For the real-world scenes in the CADAR-50K dataset, we collected additional samples using a Meta Quest 3 headset and a cellphone camera.
+The video dataset contains both malicious cognitive attacks and benign augmented reality content.
+All data was recorded using the built-in pass-through camera of a Meta Quest headset and a Samsung Galaxy S24.
+We employed two distinct data generation pipelines to ensure diversity in environmental contexts and lighting conditions.
+B.2 In-Situ Augmented Reality Overlay
+The first method involved augmenting real-world environments with manipulated overlays to simulate attacks in situ.
+We first identified target objects within a physical scene (e.g., a store sign displaying a discount).
+A high-resolution photograph of the target was captured and subjected to digital manipulation.
+For instance, altering text from ”50% off” to ”150% off” to simulate a text manipulation attack.
+The manipulated asset was imported into the Unity engine and mapped onto a planar mesh (Quad).
+We utilized Unity’s scaling and transformation tools to align the virtual overlay with the physical object’s dimensions and perspective in the real world.
+To introduce benign confounders, we analyze the scene context and suggest appropriate non-malicious objects.
+Based on these suggestions, synthetic 3D or 2D assets were generated and anchored within the scene.
+During the recording phase, a controller input triggered the visibility of the overlay, allowing us to capture the transition between the benign state and the attack state dynamically.
+B.3 Synthetic Scene Projection
+To expand the dataset beyond accessible physical locations, we utilized a fully synthetic generation pipeline.
+In the Unity environment, we instantiated a large-scale Quad (approximately $5m\times 5m$) to serve as a dynamic backdrop.
+We utilized a generative image model to synthesize realistic background scenes.
+The generation of scenes containing specific target objects is required for occlusion or manipulation attacks.
+Once the generated scene was applied to the backdrop, attack artifacts were positioned in the foreground.
+Although the background was two-dimensional, we achieved a pseudo-three-dimensional effect by strategically angling the foreground objects and the camera perspective.
+Similar to the in-situ method, the attack artifacts were toggled via controller input during recording to capture paired data samples.
 Figure 9: (a) First-person view from the AR headset: the virtual object (e.g.
@@
 (b) Third-person view showing the data collector wearing the Meta Quest 3 headset and interacting with the AR environment using handheld controllers.
-Appendix C Model Architectures and Training
-For our supervised-learning baselines—ResNet and ViViT.
+Appendix C Video Baseline Implementation Details
+We provide the full implementation details of the supervised video baselines used in our experiments: the 3D–ResNet model [hara2017learningspatiotemporalfeatures3d] and the transformer-based ViViT model [arnab2021vivitvideovisiontransformer].
+Both baselines follow standard practice in video classification and are trained under consistent data splits, preprocessing steps, and evaluation metrics.
+The descriptions below summarize the implementation in our released training code.
+C.1 Dataset Splits and Sampling
+All video models use the same dataset partitions:
+  • Train / Validation / Test: 60% / 20% / 20%, stratified by attack type and scene category.
+  • One clip is extracted per video for all splits, matching the single-view evaluation protocol.
+To mitigate class imbalance during training, we adopt a class-balanced sampling strategy. Let $c_{i}$ be the number of training videos for class $i$; class weights are:
+$w_{i}=\frac{1}{c_{i}+10^{-6}},\qquad p_{i}=\frac{w_{i}}{\sum_{j}w_{j}},$ (26)
+and training videos are sampled according to $p_{i}$. This sampling procedure matches the implementation in our 3D–ResNet code.
+C.2 Clip Construction and Preprocessing
+For both ViViT and 3D–ResNet, AR videos are processed into short clips using a uniform temporal sampling pipeline:
+  • 20-frame clips:
+    obtained usingUniformTemporalSubsample(20).
+  • Spatial resizing: frames are resized to 400 $\times$ 400.
+  • Pixel normalization: mean [0.45,0.45, 0.45], std [0.225, 0.225, 0.225].
+Frames are arranged as tensors of shape $(T,C,H,W)$ with $T=20$. No additional augmentations (cropping, flipping, jittering) are applied, avoiding synthetic distortions that could alter AR semantics.
+C.3 3D–ResNet Baseline
+Our 3D CNN baseline is implemented using the r3d_18 architecture from torchvision (18-layer 3D–ResNet).We detail the full setup below.
+Architecture.
+  • Model: R3D-18 (3D ResNet).
+  • Pretraining: Kinetics-400 weights from Torchvision.
+  • Final classifier layer replaced with a 5-way linear head.
+Loss Function.
+We employ a Focal Loss with $\gamma=2.0$:
+$\mathcal{L}_{\mathrm{focal}}=(1-\exp(-\mathrm{CE}(\mathbf{x},y)))^{\gamma}\,\mathrm{CE}(\mathbf{x},y),$ (27)
+which improves stability under severe class imbalance.
+Optimization.
+  • Optimizer: AdamW
+[…]
+[…]
+  • Scheduler: StepLR (step=10, $\gamma$=0.1)
+  • Batch size: 20
+  • Epochs: 20
+Precision Training.
+Mixed-precision (FP16) training is performed using HuggingFace Accelerate with autocast() enabled, matching our code.All layers are trainable; no freezing is applied.
+C.4 ViViT Baseline
+We adopt the ViViT-B/16x2 (Factorised Encoder) model as our transformer baseline:
+  • Patch size: $\times16$
+  • Hidden dimension: 768
+  • Transformer layers: 12
+  • Heads: 12
+  • Pretraining: Kinetics-400
+The model receives the same 20-frame clips and performs temporal factorization before spatial tokenization as in the original ViViT design.
+Training.
+We train ViViT for 20 epochs using:
+  • Batch size: 16
+  • AdamW optimizer (1e-4 learning rate, 5e-3 weight decay)
+  • Cosine learning-rate decay
+  • Warmup steps: 10% total steps
+  • Mixed precision (FP16) using Accelerate
+C.5 Evaluation Metrics
+To ensure consistency with the metrics defined in the main paper, we evaluate both video models using the following measures:
+Class-Specific Accuracy.
+For each class $k$:
+$\text{Acc}_{k}=\frac{\mathrm{TP}_{k}}{\mathrm{TP}_{k}+\mathrm{FN}_{k}+\mathrm{FP}_{k}}.$
+Precision, Recall, F1.
+Computed using scikit-learn with zero_division=0.
+Confusion Matrix.
+Collected over the test split and used to compute TP/FP/FN counts for each class.
+C.6 Fairness and Consistency
+To ensure a fair comparison with CADAR and VLM baselines:
+  • All models receive the same raw input videos.
+  • No prompts, symbolic structures, or additional annotations are provided to video baselines.
+  • Clip preprocessing (20 frames, resizing, normalization) is identical across all supervised baselines.
+Overall, these settings yield strong and transparent supervised video baselines, allowing us to isolate the contributions of CADAR’s neurosymbolic reasoning.
+C.7 Model Architectures and Training
+For our supervised-learning baselines—3D ResNet and ViViT.
 Training hyper-parameters are shown below.
@@
 batch size 20 [8–20] 3 16 [4–16] 2
-epochs 10 [5–50] 6 10 [5–50] 4
+epochs 20 [5–50] 6 20 [5–50] 4
 scheduler StepLR Cosine
@@
 All input videos are resized to 400×400 pixels and normalized using a mean of [0.45, 0.45, 0.45] and a standard deviation of [0.225, 0.225, 0.225].
-Each input consists of eight uniformly sampled frames from a 13-second clip.
+Each input consists of 20 uniformly sampled frames from a video clip.
 We use Focal Loss with a gamma of 2.0 as the training objective.
@@
 Final hyperparameter configurations for both models were selected based on the settings that yielded the highest validation accuracy.
-Appendix D Benchmark Prompt Templates (GPT-o3 and Gemini 2.5)
-For our baseline VLM models, we use the following prompt for both GPT-o3 and Google Gemini 2.5.
-Related work include(Zhang et al.
-2025; Zou et al.
-2024; Qiao et al.
-2024; Zhang et al.
-2024b; Jing et al.
-2024; Gao et al.
-2024; Ravari et al.
-2024; Zhang, Aggarwal, and Lan 2025; Yu et al.
-; Zhang, Imani, and Lan 2024)
+Appendix D Vision–Language Model (VLM) Baseline Implementation Details
+We evaluate three state-of-the-art multimodal VLMs as baselines: GPT-5-mini (OpenAI), Gemini-2.5-flash (Google), and Grok-4-fast (xAI).
+All three models support multi-image inputs and natural language reasoning, enabling them to perform frame-level attack classification without any training or fine-tuning.
+This section details our prompt design, image preprocessing strategy, and inference procedure, following our released implementation.
+D.1 Input Frames and Preprocessing
+Each AR video is converted into multiple reference and attack frames using the procedure in our VLM baseline code.We extract:
+  • N reference frames uniformly from the first 30–40% of the video, and
+  • M attack frames uniformly from the last 25%.
+Images are resized to a maximum resolution of $800\times 800$ while preserving aspect ratio.
+For GPT-5 and Grok, images are encoded as JPEG base64 strings.
+Gemini receives PIL image objects directly, as required by the generate_content interface.
+D.2 Unified Prompt for All VLM Baselines
+To ensure fair comparison across models, all VLMs use the exact same prompt (text only; no model-specific tuning).
+We introduce the four base attack types used in CADAR—text modification, visual modification, obstruction, and injection—together with detailed natural-language definitions.
+The full prompt is shown below.
 You are an AR video attack analyst.
-Your job is to identify the presence of 5 different kinds of AR attacks in the video.
-The definitions for each attack are as seen below: A text modification attack manipulates textual information embedded in the environment, thereby changing the semantic message conveyed by the scene.
-For example, replacing a NO PARKING FIRE LANE sign with one reading FREE PARKING FIRE LANE, for instance, directly reverses the intended directive and can create significant safety hazards.
-Such attacks are especially consequential in contexts where text offers primary contextual or navigational cues.
-A visual modification attack alters an object’s appearance—its color, shape, or iconography—or repositions it in an implausible manner.
-Appearance changes, such as transforming a pedestrian-crossing sign into a bicycle-crossing sign or switching a green traffic signal to red, degrade recognition accuracy and mislead both human users and perception systems.
-Spatial changes include shifting an object away from its logical location or removing it entirely; for example, relocating a stop sign to the centre of an intersection.
+Your job is to identify the presence of 4 different kinds of AR attacks and non-attacks in the video with the scene context.
+The definitions for each attack are as follows:
+A text modification attack manipulates textual information embedded in the environment, thereby changing the semantic message conveyed by the scene.
+For example, replacing a “NO PARKING FIRE LANE” sign with “FREE PARKING FIRE LANE” directly reverses the intended directive and can create safety hazards.
+Such attacks are consequential when text provides primary contextual or navigational cues.
+A visual modification attack alters an object’s appearance—its color, shape, or iconography—or repositions it implausibly.
+Examples include transforming a pedestrian-crossing sign into a bicycle-crossing sign, switching a green traffic signal to red, or relocating a stop sign to the center of an intersection.
 These manipulations can be quantified by analysing pixel-level discrepancies and deviations from expected spatial distributions.
-A removal attack hides or deletes critical objects, thereby eliminating essential information and distorting scene semantics.
-In a fire evacuation scenario, a blank patch that obscures an emergency-exit sign on a door.
-By erasing this safety-critical cue, the attack deprives occupants of clear evacuation guidance, potentially endangering lives during an emergency.
-By analyzing nodes from the scene graph, removal attacks disrupt relational structure and semantic integrity.
-An addition attack introduces new, contextually inconsistent, or misleading objects into the scene, thereby distorting its intended semantics or altering user perception.
-For example, a fake road sign indicating a detour is inserted into a clear street, potentially redirecting traffic and causing confusion.
-These attacks may include adding fictitious warning signs, virtual obstructions, or irrelevant icons that do not exist in the physical environment.
-Such manipulations exploit the overlay nature of AR systems and are particularly dangerous when added elements resemble legitimate cues, misleading both human users and perception models.
+An obstruction attack hides or deletes critical objects, thereby eliminating essential information and distorting scene semantics.
+For example, erasing an emergency-exit sign in a fire evacuation scenario removes a safety-critical cue and disrupts the relational structure of the environment.
+An Injection attack introduces new, contextually inconsistent, or misleading objects that distort scene semantics.
+A fake detour sign placed on a clear street, or fictitious warning icons inserted into an AR overlay, may mislead human users and perception systems.
 Detection relies on identifying unexpected objects that conflict with the scene’s established structure and context.
-The following input will contain a video.
-Please output the attack kind.
-Note that there may be more than 1 kind of attack present in the video, or a video could have no attack present at all.
-Along with the attack, mention the object the attack was performed on or describe the attack briefly in the context of the video.
+A non-attack refers to a video in which the AR content remains semantically consistent with the physical environment and does not alter the user’s understanding of the scene.
+The following input will contain video frames and context.
+Please output the class.
+Note that there may be more than one kind of attack present in the video, or no attack at all.
+Along with the attack, mention the object the attack was performed on, or briefly describe the attack in the context of the video.
+This prompt is prepended to the multimodal input for all VLMs.
+D.3 Inference Procedure
+Each VLM receives:
+  • the textual prompt above,
+  • scene context,
+  • $N=20$ reference images, and
+  • $M=5$ attack image at a time.
+The models are called with their native multimodal APIs:
+  • GPT-5-mini: OpenAI Chat Completions (image URLs encoded as base64).
+  • Gemini-2.5-flash: generate_content with a mixed list of text + images.
+  • Grok-4-fast: xAI Chat Completions (OpenAI-compatible JSON format with image URLs).
+Each response is required to be valid JSON, following the schema below:
+    "object": "<object>",
+    "attack_type": "<classification>",
+    "notes": "<short explanation>"
+We apply a lightweight parser that extracts JSON blocks from the raw VLM outputs.
+Each model produces one prediction per attack frame.
+This mirrors video-level behavior and reduces variance inherent in single-frame VLM outputs.
+D.4 Fairness Measures
+To ensure a strict apples-to-apples comparison:
+  • All VLMs use the same prompt with context, the same set of reference frames, and the same attack frames.
+  • No model-specific prompt engineering, priming, or handcrafted reasoning chains are used.
+  • All text is fed to each VLM exactly once per attack frame; no iterative refinement or self-consistency sampling is applied.
+Because VLMs are not adapted to CADAR’s symbolic representation or temporal structure, these baselines reflect the true performance of general-purpose open-world vision–language models on cognitive attack detection.